### v1

# Learning Priors for Non Rigid SfM from Casual Videos

Abstract
We tackle the long-standing challenge of reconstructing 3D structures and camera positions from videos. The problem is particularly hard when objects are transformed in a non-rigid way. Current approaches to this problem make unrealistic assumptions or require a long optimization time. We present TracksTo4D, a novel deep learning-based approach that enables inferring 3D structure and camera positions from dynamic content originating from in-the-wild videos using a single feed-forward pass on a sparse point track matrix. To achieve this, we leverage recent advances in 2D point tracking and design an equivariant neural architecture tailored for directly processing 2D point tracks by leveraging their symmetries. TracksTo4D is trained on a dataset of in-the-wild videos utilizing only the 2D point tracks extracted from the videos, without any 3D supervision. Our experiments demonstrate that TracksTo4D generalizes well to unseen videos of unseen semantic categories at inference time, producing equivalent results to state-of-the-art methods while significantly reducing the runtime compared to other baselines. Project page: https://tracks-to-4d.github.io.
Keywords:
Structure From Motion Dynamic Videos 3D Reconstruction Equivariance Symmetries
Figure 1: We present TracksTo4D, a method for mapping a set of 2D point tracks extracted from casual dynamic videos into their corresponding 3D locations and camera motion. At inference time, our network predicts the non-rigid structure and camera motion in a single feed-forward pass. Top. An example showing the input video frames ordered from left to right. A set of pre-extracted point tracks that are used as input to our network and presented in corresponding colors. Bottom. The outputs of our network for 3 time steps as observed from the right side of the scene. The camera trajectory is present as gray frustums, whereas the current camera is marked in red. The reconstructed 3D scene points are presented in corresponding colors to the input tracks on the top. See video visualizations here.
1 Introduction
Predicting 3D structures in dynamic scenes is a challenging problem. In contrast to rigid scenes where the epipolar geometry constraints hold between the corresponding points of different views [17], determining the depth of a moving point from monocular views is an ill-posed problem. Moreover, while multiview camera pose estimation can be solved successfully by Structure from Motion methods [37, 53, 30], these techniques tend to fail in scenarios with significant non-rigid motion [21].
Background and previous work.
Unfortunately, current approaches to solving this problem are based on unrealistic assumptions or suffer from long optimization times per scene that make them inapplicable to real-world problems. More concretely, early techniques for Non-Rigid SFM [6] used matrix factorization of long-term point tracks matrix of the non-rigid scene part. This was done by assuming that a small linear basis and per-frame coefficients can formulate the dynamic structure. By further assuming an orthographic camera is used, they omitted camera translations and solved for rotations and 3D structure by matrix factorization. While follow-up work improved accuracies [22, 11], this setup is still not practical for real-world casual videos, captured by pinhole cameras. Recently, the Neural Radiance Field (NeRF) [28] concept has brought major progress to novel view synthesis and surface reconstruction [60, 49] for static scenes. While some follow-up work extended NeRF to non-rigid scenes [23, 24, 33], optimizing a non-rigid NeRF requires a long, and per-scene optimization. Other recent approaches [26, 21, 62, 63] optimized for consistent depth maps given a specific monocular video by fine-tuning pre-trained single-view-depth-predictions [34]. These methods showed impressive progress in camera and non-rigid structure estimation, but also require per-video optimization at inference time. Other approaches learned prior and parameterizations for specific non-rigid object classes such as humans [36, 25]. This prior enables shorter inference times for obtaining 3D structures from monocular videos that contain instances from the learned category [55]. However, this approach does not generalize well across shape categories and mostly requires per-video finetuning.
Our approach.
In this work, we propose TracksTo4D, a novel deep learning-based approach for predicting the 3D dynamic Non-Rigid Structure And camera Motion from videos. Our key observation is that the 2D motion present in casual videos shares common patterns across a variety of semantic video categories. Crucially, and as was suggested by [6], this motion can be effectively encoded into a point track matrix, where each entry represents the 2D position of a tracked point in a specific video frame. Rather than processing image pixels or semantic features that vary dramatically across different scenes and objects, we propose directly processing these point track matrices to learn generic, class-agnostic features. This also differs from the common approach of using optical flow representations of motion, as utilizing optical flow fields as input to deep neural networks is non-trivial.
TracksTo4D  builds on the recent progress in point tracking [12, 18], which introduced the ability to extract long-term point tracks from casual videos. We leverage this progress and design a deep neural network architecture, that maps the point tracks of a given video into corresponding camera poses, 3D points, and per-point non-rigid motion level values (See Fig. 1). Crucially, the mapping we aim to learn preserves two natural symmetries: (1) the points being tracked can be arbitrarily permuted without affecting the problem; (2) the frames containing these points exhibit temporal structure, adhering to an approximate time-translation symmetry. Following the Geometric Deep Learning paradigm [7], we build upon recent theoretical advances in equivariant learning [27] and integrate these two symmetries into our network architecture using dedicated attention mechanisms.
Predicting the camera poses and the 3D non-rigid motion from 2D motion alone is an ill-posed problem without any constraints, and equivalent to finding the 3D structure from each frame [3]. Inspired by previous work [6], we model the predicted 3D point cloud as a weighted sum of a 3D structure basis, where the first basis is assumed to fully represent the rigid points, and the rest of the basis elements are used to complete the non-rigid deviation for the non-rigid points. The predicted motion level values determine the non-rigidity of each 3D world point, i.e. the inconsistency of the rigid assumption that is represented by the first basis element for this world point.
In our experiments, TracksTo4D is trained on the Common Pets dataset [38] using only 2D point tracks extracted by [18], without any 3D supervision by simply minimizing the reprojection errors. Interestingly, we observe that by just "watching" point tracks in videos, our network learns to infer their 3D location and the camera motion. We evaluate our method on test data with GT depth maps and demonstrate that it generalizes well, and produces comparable results to the state-of-the-art methods, with significantly shorter inference time.
Contributions.
In summary, our contributions are: (1) Designing an equivariant network architecture that maps from 2D point trajectories into corresponding non-rigid structures and camera poses; (2) Introducing a training paradigm to identify the rigid part of the scene for constraining the camera poses; (3) Simple training on casual videos without any 3D supervision; (4) Demonstrating fast inference time, accurate results, and strong generalization across animal categories.
2 Related Work
Structure from Motion
The Structure from Motion (SfM) problem recovers the 3D structure and the camera poses from images. Early methods [44, 39] factorized a 2D measurement matrix of corresponding image points into a multiplication of a camera poses matrix by a 3D points matrix. These methods assume that each recovered 3D point is tracked over the entire image sequence which is not the case in practical use cases. Incremental approaches for SfM [37, 1, 53] start from a minimal set of images, to build an initial reconstruction. Then, the rest of the images are added to the scene one by one, by finding 2D correspondences from the new image into scene points. To prevent drift, Bundle Adjustment [45] is repeatedly applied to the reconstruction during this process. Several global methods [52, 10, 19, 51, 32] reduced the running time by using relative two-view geometries between image pairs to extract absolute camera poses in a single step, followed by a final round of Bundle Adjustment.
Recently, Moran et al. [29] presented a method for applying an order equivariant architecture for mapping a set of multiview image correspondences into two sets of cameras and 3D points. It was trained on a dataset of scenes and obtained a reduced optimization time on novel scenes, due to the learned prior. A follow-up work [8] improved generalization by adding graph attention layers into the architecture of Moran et al. Other recent deep learning-based techniques used differentiable Bundle Adjustment layers [47, 40], or diffusion models [61, 48]. The above-mentioned techniques assumed that the structure is rigid such that the whole scene can be represented in a single 3D structure.
Simultaneous Localization and Mapping (SLAM)
In contrast to SfM which assumes unordered images, monocular Simultaneous Localization and Mapping (SLAM) methods [30, 31, 15] get as input a sequence of video frames recorded by a single camera. These methods often define a set of keyframes for reconstructing the scene with a mapping thread, while registering every new frame into the structure by a localization thread. The mapping thread repeatedly applies Bundle Adjustment for updating the structure. ORB-SLAM [30] uses feature points for mapping a sparse scene representation and tracks the camera with sparse point matches. Other methods [31, 15, 14] use denser scene representations and align cameras directly by using photometric constraints.
More recent methods combine classical approaches with deep learning-based tools e.g. depth prediction from a single image, optical flow prediction, or relative camera pose estimation [59, 5, 50, 65, 66, 41]. As in SfM, these methods assume rigid scenes, fail to produce the cameras in scenes with large portions of non-rigid motion, and cannot reproduce non-rigid parts of the scene. DROID-SLAM [43] used synthetic data with ground truth 3D supervision for learning to predict camera poses via deep-based bundle adjustment on keyframes while excluding dynamic objects. ParticleSfM [64] pre-trained a motion prediction network on a synthetic dataset with ground truth supervision, and used it to filter out any non-rigid content for reproducing the cameras in dynamic scenes. Both, [43, 64] do not infer the non-rigid 3D structure.
Test-Time Optimization for Non-Rigid Scenes
Recent methods [26, 21, 62, 63] finetuned the monocular depth estimation from a pre-trained model [35, 34] using optical flow constraints [42], for obtaining consistent dense depth maps for a monocular video. [62] further optimized motion maps that represent the $\gamma$ parameters of zero mean Cauchy distributions for handling scenes with large dynamic motion. [54, 16] use depth from single image estimations, to improve novel view synthesis in dynamic scenes. [24] further optimized for the unknown camera poses together with the dynamic radiance field optimization. LASR [56], ViSER [57] and BANMo [58] optimize for a dynamic surface given a monocular video by assuming rigid bones and linear blend skinning weights. However, all the above-mentioned methods require per-scene optimization, resulting in slow inference. Recently, [38] presented the Common Pets in 3D (COP3D) dataset that contains casual, in-the-wild videos of pets, and used it to learn priors for novel view synthesis in dynamic scenes.
3 Problem formulation
Our problem setup is illustrated in Fig. 2. Given a video of $N$ frames, let $M\in\mathbb{R}^{N\times P\times 3}$ be a pre-extracted point tracks tensor (Fig. 2, left side). This tensor represents the two-dimensional information about a set of $P$ world points that are tracked throughout the video. Each element in the tensor, $M_{i,j}$, stores three values: $(x,y,o)$ where $x,y\in\mathbb{R}$ are respectively the observed horizontal and vertical locations of point $j$ in frame $i$, and $o\in\{0,1\}$ indicates whether point $j$ is observed in frame $i$ or not. Our goal is to learn a function $f$ to map the input point tracks tensor $M$ into a set of per-frame camera poses $\{R_{i},\mathbf{t}_{i}\}_{i=1}^{N}$ and per-frame 3D points $\{X_{i}\}_{i=1}^{N}$, where $R_{i}\in SO(3),\mathbf{t}_{i}\in\mathbb{R}^{3},X_{i}\in\mathbb{R}^{P\times 3}$ (Fig. 2, right side).
Discussion.
We note that given enough point tracks, $M$ contains the 2D motion constraints that can be provided by the video. The source of each 2D motion constraint can be either from a global camera 3D motion or from an independent object motion. While the pixel appearance and semantic information could be used as well as inputs in our task, we find that the 2D motion signal $M$ gives useful geometric constraints for our setup. Moreover, this abstraction of the video helps our network generalize well to out-of-distribution videos with object classes that have not been observed during training but share similar 2D motion characteristics as we show in our experiments. Therefore $f$ only uses $M$ as input and does not rely on the input video once $M$ has been extracted.
Figure 2: Our pipeline Our network takes as input a set of 2D point tracks (left) and uses several multi-head attention layers while alternating between the time dimension and the track dimension (middle). The network predicts cameras, per-frame 3D points, and per-world point movement value (right). The 3D point colors illustrate the movement level values, such that points with high/low motion values are presented in blue/purple colors respectively. These outputs are used to reproject the predicted points into the frames for calculating the reprojection error losses. See details in the text.
4 Method
4.1 Overview
Our pipeline is illustrated in Fig. 2. Our method receives $M\in\mathbb{R}^{N\times P\times 3}$ as input. This tensor is being processed by a neural architecture composed of multi-head attention layers where the attention is calculated sequentially on the $P$ and the $N$ dimensions in each layer (Sec. 4.3). After several such layers, we perform a projection to $N$ per-frame parameters and $P$ per-world-point parameters which determine the camera poses and 3D dynamic structure (Sec. 4.2). This projection is done by first pooling over the secondary dimension and applying suitable layers to each one of the outputs to get the per-camera and per-world points parameters respectively. Our network is trained by minimizing the reprojection error and other regularization losses (Sec. 4.4) that are used to update the model parameters.
4.2 Constraining the Non-Rigid Structure and Motion
Given the input 2D tracks, we wish to describe their motion by global camera motion and non-rigid self-motion. The 2D motion that is produced from rigid points constrains well the camera motion, but predicting the camera poses and the 3D non-rigid motion from 2D motion alone is an ill-posed problem without any constraints [3].
We tackle this challenge in two ways. First, as done in classical Non-Rigid Structure from Motion [6], we constrain the output points to be formulated by a linear combination of basis elements. This constrains the 3D motion to be simple and prevents "jumps" in the unconstrained depth. More formally we assume that the non-rigid structure is formulated by a few basis elements $B_{1},\dots,B_{K}\in\mathbb{R}^{P\times 3}$ such that $X_{i}=\sum_{k=1}^{K}c_{ik}B_{k}$ where $\{c_{ik}\}_{k=1}^{K}$ are the $K$ coefficients which define the 3D structure of frame $i$.
Secondly, we define the first basis element $B_{1}\in\mathbb{R}^{P\times 3}$ to be a rigid approximation of the points, such that the first basis coefficient in every frame is $1$, i.e. $c_{i1}=1$ for $i=1,\dots,N$. This helps in constraining the camera motion from the 2D motion of the rigid scene part. However, since this global rigid approximation cannot explain the 2D motion that is produced by the non-rigid world points, we model the reprojection error of the rigid approximation with $P$ zero-mean Cauchy distributions. For a non-rigid world point, $B_{1}$ by itself cannot produce low reprojection errors, and therefore its associated $\gamma$ is expected to be higher. The parameters of the Cauchy distributions, $\gamma_{1},\dots\gamma_{P}\in\mathbb{R}_{+}$ (Fig. 2, "Movement Level Values") are predicted as well by the network and our training maximizes the likelihood of this assumption (see details in Sec. 4.4). This allows our network to benefit from the rigid part for camera pose prediction while still being robust to the non-rigid motion. The rest of the basis elements determine the non-rigid deviation of the non-rigid points from the rigid approximation.
In summary, our function $f$ gets $M\in\mathbb{R}^{N\times P\times 3}$ as an input and predicts per-frame outputs:
$R_{i}\in SO(3),\mathbf{t}_{i}\in\mathbb{R}^{3},\{c_{ik}\}_{k=2}^{K},\text{for}\ \ i=1,\dots,N$ (1)
and per-point outputs:
$B_{1},\dots,B_{K}\in\mathbb{R}^{P\times 3},\gamma_{1},\dots,\gamma_{P}\in\mathbb{R}_{+}$ (2)
which together define the non-rigid structure and motion.
4.3 Architecture
Following the geometric deep learning paradigm, our goal is to design a neural architecture that respects the underlying symmetries and structure of the data.
Symmetry analysis.
Our input is a tensor $M\in\mathbb{R}^{N\times P\times 3}$ representing a sequence of $N$ frames each with $P$ point coordinates. This structure gives rise to two key symmetries: First, the order of the $P$ points within each frame does not matter - permuting this axis results in an equivalent problem [27]. Formally, this axis has a permutation symmetry $S_{P}$ where $S_{P}$ is the symmetric group on $P$ elements. Second, along the temporal $N$ axis, we have an approximate translation symmetry arising from the ordered video sequence. This means that translating the time frames is required to result in nearly the same overall representation. We model this with a cyclic group $C_{N}$ of order $N$. Both symmetries are illustrated in Fig. 3.
Figure 3: The symmetry structure of our problem. Frames (vertical) have time translation symmetry while points (horizontal) have set permutation symmetry.
Taken together, the full symmetry group of the input space is the direct product $\mathcal{G}=C_{N}\times S_{P}$ combining these frame and point permutation symmetries  [^1] This is different from the symmetry group studied in [29], where the temporal structure was not exploited. Next, we will design an architecture equivariant to $\mathcal{G}$, so that we can incorporate these domain priors and ensure that the model takes into account the underlying data structure.
Network Architecture.
Building on the analysis from the Deep Sets for Symmetric elements (DSS) approach [27], if we assumed we can use only linear layers, our architecture would take the form:
$F(M)_{:,j}=L_{1}(M_{:,j})+\sum_{j^{\prime}=1}^{P}L_{2}(M_{:,j^{\prime}})$ (3)
where $L$ is a translation equivariant function, $M_{:,j}\in\mathbb{R}^{N\times d}$ are the columns of $M$ representing all the inputs for a specific tracked point, $F(M)_{:,j}\in\mathbb{R}^{N\times d^{\prime}}$ is the output column and we have $d,d^{\prime}$ input and output feature channels respectively. While valid, this linear layer design might not be the optimal choice. To enhance the model, we incorporate additional nonlinear layers that are well-suited for sequence and set structures data. One particularly effective choice is the transformer layer [46]. Importantly, by leveraging attention mechanisms and temporal positional encodings, the transformer processes sequential input while respecting its ordering and context. Specifically, our layer $F$ is formulated very similarly to Equation (3) and is implemented by replacing the convolutions ($L_{i}$) and summations in Equation (3) with two self-attention mechanisms and suitable temporal positional encoding across the $N$ dimension. We provide more architecture and implementation details in the supplementary.
4.4 Training
Given a dataset of point track representations: $\{M_{1},\dots,M_{L}\}$, our method trains the function $f$ to minimize the loss function over all elements in the dataset:
[…]
Our loss function is defined by:
$\mathcal{L}=\lambda_{\text{Reproject}}\mathcal{L}_{\text{Reproject}}+\lambda_{\text{Rigid}}\mathcal{L}_{\text{Rigid}}+\lambda_{\text{Negative}}\mathcal{L}_{\text{Negative}}+\lambda_{\text{Sparse}}\mathcal{L}_{\text{Sparse}}$ (5)
The first loss is the reprojection error given by the structure and the cameras predicted by $f$:
$\mathcal{L}_{\text{Reproject}}=\frac{1}{\sum_{i=1}^{N}\sum_{j=1}^{P}M_{ij}^{o}}\sum_{i=1}^{N}\sum_{j=1}^{P}M_{ij}^{o}\mathcal{R}(X_{ij},R_{i},\mathbf{t}_{i},M_{ij}^{xy})$ (6)
where $\mathcal{R}(X_{ij},R_{i},\mathbf{t}_{i},M_{ij}^{xy})$ is the reprojection error when projecting the point $X_{ij}$ with the camera pose $(R_{i},\mathbf{t}_{i})$ with respect to the measured point $M_{ij}^{xy}$:
$\mathcal{R}(X_{ij},R_{i},\mathbf{t}_{i},M_{ij}^{xy})=\left\lVert\frac{(R_{i}^{T}(\mathbf{X}_{ij}-\mathbf{t}_{i}))_{1,2}}{(R_{i}^{T}(\mathbf{X}_{ij}-\mathbf{t}_{i}))_{3}}-M_{ij}^{xy}\right\rVert$ (7)
As discussed in Sec. 4.2, the first predicted basis element $B_{1}\in\mathbb{R}^{P\times 3}$ defines a rigid (fixed in time) approximation. Our network also predicts a movement coefficient $\gamma_{j}$ for each world point that defines a zero-mean Cauchy distribution. Given $\gamma_{j}$ and the reprojection error $r_{ij}=\mathcal{R}(B_{1j},R_{i},\mathbf{t}_{i},M_{ij}^{xy})$ of the $j^{th}$ point of $B_{1}$ that is projected by the $i^{th}$ camera, the negative log-likelihood of $r_{ij}$ distributed according to the $\gamma_{j}$-zero-mean Cauchy distribution is proportional to:
$\mathcal{C}(r_{ij},\gamma_{j})=\log\left(\gamma_{j}+\frac{r_{ij}^{2}}{\gamma_{j}}\right)$ (8)
Our rigid loss calculates the mean negative log-likelihood over all observed points in all frames:
$\mathcal{L}_{\text{Rigid}}=\frac{1}{\sum_{i=1}^{N}\sum_{j=1}^{P}M_{ij}^{o}}\sum_{i=1}^{N}\sum_{j=1}^{P}M_{ij}^{o}\mathcal{C}(\mathcal{R}(B_{1j},R_{i},\mathbf{t}_{i},M_{ij}^{xy}),\gamma_{j})$ (9)
As in [29] we regularize the points to be in front of the camera by:
$\mathcal{L}_{\text{Negative}}=-\sum_{i=1}^{N}\sum_{j=1}^{P}M_{ij}^{o}\text{ Min}(R_{i}^{T}(\mathbf{X}_{ij}-\mathbf{t}_{i}))_{3},0)$ (10)
We further find it beneficial to regularize the deviation from the rigid approximation $B_{1}$ to be sparse at locations with low $\gamma$ values:
$\mathcal{L}_{\text{Sparse}}=\frac{1}{P(K-1)}\sum_{k=2}^{K}\sum_{j=1}^{P}\frac{1}{3\gamma_{j}}\left(|B_{kj1}|+|B_{kj2}|+|B_{kj3}|\right)$ (11)
where $\gamma$ is detached from the gradient calculation for this loss.
ATE $\downarrow$ RPE Trans $\downarrow$ RPE Rot $\downarrow$ Inference $\downarrow$
(mm) (mm) (deg) Time
DROID-SLAM [43] 5.08 3.60 0.20 10 seconds
ParticleSfM [64] 12.79 6.95 0.51 11 minutes
RCVD[21] 43.95 25.77 2.31 20 minutes
CasualSAM [62] 6.90 3.95 0.22 130 minutes
Ours (cats only) 8.96 3.79 0.23 0.16(+8.6) seconds
Ours (cats only)+BA 4.22 2.86 0.17 0.4(+8.6) seconds
Ours (dogs only) 8.03 3.54 0.23 0.16(+8.6) seconds
Ours (dogs only)+BA 4.19 2.83 0.17 0.4(+8.6) seconds
Ours (dogs+cats) 8.11 3.68 0.24 0.16(+8.6) seconds
Ours (dogs+cats)+BA 4.21 2.86 0.17 0.4(+8.6) seconds
Table 1: Camera Poses Accuracy. Quantitative comparisons between our method to previous methods for camera accuracy and running time. For our method, we compare three versions of (1) training only on cats, (2) training only on dogs, and (3) training on both. We evaluate all methods on an unseen test set of cats and dogs. We further show our accuracies after one round of Bundle Adjustment. For our method’s running time, we also mention the point tracks extraction time in parenthesis (+8.6 seconds) which is performed by [18] as a preprocess.
Abs Rel $\downarrow$ $\delta<1.25\uparrow$ $\delta<1.25^{2}\uparrow$ $\delta<1.25^{3}\uparrow$
Dynamic All Dynamic All Dynamic All Dynamic All
RCVD[21] 0.40 3.6e+7 0.43 0.72 0.75 0.90 0.92 0.96
MiDaS [4] 0.16 6.1e+4 0.78 0.71 0.97 0.88 1.00 0.93
CasualSAM [62] 0.09 0.06 0.93 0.97 0.99 0.99 1.00 1.00
Our (cats only) 0.11 0.08 0.88 0.92 0.99 0.98 1.00 1.00
Our (dogs only) 0.12 0.08 0.85 0.91 0.99 0.99 1.00 1.00
Our (dogs+cats) 0.12 0.08 0.85 0.91 0.98 0.98 1.00 1.00
Table 2: Depth Accuracy We show a comparison to previous methods on the predicted depth for the trajectories compared to their GT depths. We compare three instances of our method, each trained on a different partition of the dataset.
Abs Rel $\downarrow$ $\delta<1.25\uparrow$ $\delta<1.25^{2}\uparrow$ $\delta<1.25^{3}\uparrow$ Rep.(pix.) $\downarrow$ ATE $\downarrow$ RPE Trans $\downarrow$ RPE Rot $\downarrow$
Dyn. All Dyn. All Dyn. All Dyn. All Dyn. All (mm) (mm) (deg)
Set of Sets 0.27 0.15 0.60 0.77 0.87 0.94 0.97 0.99 9.86 5.33 16.87 5.53 0.39
No $\mathcal{L}_{\text{Rigid}}$ 0.77 0.36 0.25 0.46 0.48 0.70 0.68 0.82 1.00 0.86 96.20 29.86 0.99
No $\gamma$ 0.22 0.16 0.66 0.73 0.93 0.91 0.99 0.97 4.54 2.41 13.91 4.86 0.29
K=30 0.14 0.09 0.81 0.90 0.97 0.98 0.99 0.99 4.88 2.78 9.39 3.68 0.23
K=2 0.11 0.08 0.88 0.91 0.98 0.98 1.00 1.00 8.58 3.56 9.31 3.86 0.25
DSS 1.65 0.58 0.19 0.35 0.34 0.60 0.47 0.74 63.75 70.60 34.90 22.63 1.64
No $\mathcal{L}_{\text{Sparse}}$ 0.17 0.13 0.79 0.80 0.95 0.94 1.00 0.99 4.57 2.73 11.79 7.99 0.55
Full 0.11 0.08 0.88 0.92 0.99 0.98 1.00 1.00 3.98 1.97 8.96 3.79 0.23
Table 3: Ablation Study. We evaluate the contribution of different parts from our method by changing one property each time. See details in the text.
Figure 4: Qualitative Results. We present qualitative results in the same format as in Fig. 1. Top. 3 frames from 2 different test video sequences with point tracks marked with corresponding colors. Bottom. A 3D visualization of the outputs of our method, from two time stamps. As the scene is observed from the same viewpoint in each pair, a non-rigid movement can be observed.
5 Experiments
5.1 Training Details
We trained our network on the cat and dog partitions from the COP3D dataset [38], which contains a diverse set of casual real-world videos of pets. We trained our network from scratch three times to test our generalization capability between semantic categories: once on the cat partition, once on the dog partition, and once on both partitions combined. In total, we used $733$ cat videos and $753$ dog videos for training. We trained our networks for $7000$ and $3500$ epochs for the single-class and multi-class setups respectively. Training our method lasts about one week on a single Tesla V100 GPU with 32GB memory. We used the Adam optimizer [20] with a learning rate of $10^{-4}$. For our non-rigid bases $B_{1},...,B_{K}$ (Sec. 4.2) we used $K=12$. Our method assumes known camera internal parameters which are provided by the dataset and used to normalize the point tracks as a preprocess.
5.2 Evaluation Data
While the COP3D dataset provides cameras that were extracted by COLMAP [37], we note that the non-rigid structure was captured as well in part of their reconstruction which indicates that its reconstruction might not be accurate. Furthermore, the coordinates system units of these reconstructions are unknown. Finally, this dataset does not have any depth map information for evaluating the non-rigid structure.
For evaluation, we used an RGBD sensor and collected 21 casual videos of dogs and cats, each with 50 frames. The depth maps were used as ground truth for evaluating the reconstructed structure. We extracted the cameras by running COLMAP on the images while masking out the pet areas with dilatated masks provided by [68]. The cameras were scaled to millimeter units using the provided GT depth. We verified that this data includes enough non-rigid motion, by also including several videos that COLMAP failed to reconstruct without the masks. We further verified manually that the camera trajectories look reasonable. Note that this is test data that our network did not see during training and was not used to tune our hyperparameters.
5.3 Results
We evaluated our method on the test data. Qualitative visualizations are presented in Fig. 1 and Fig. 4. We also show a visualization of the movement level values, $\gamma$ in Fig. 5. We compared our method to the state-of-the-art methods and evaluated both, the camera poses and the structure. The camera poses were evaluated using the Absolute Translation Error(ATE), the Relative Translation Error(RTE), and the Relative Rotation Error(RRE) metrics after coordinates system alignment. The results are presented in Tab. 1. We compare three training possibilities of our method of training only on cats, only on dogs, and on both. As can be seen in the table, the results are similar in all 3 cases, even though the dest set contains samples of both dogs and cats. This shows that our network generalizes well to unseen animal classes. Our output camera poses as inferred by the network are already accurate and outperform some of the prior methods. We further show the results of our method after a single and short round of Bundle Adjustment, which makes our method the most accurate one.
Tab. 1 also compares the running times between the methods. It can be seen that our method, even with bundle adjustment is the fastest one when most of the time (8.6 seconds) is used to extract the input point tracks with [18].
To evaluate the structure we compare our method using depth evaluation metrics [13]. The results are presented in Tab. 2, and show that our inferred structure is comparable to the state-of-the-art [62] while taking significantly shorter running times (a few seconds for our method versus more than two hours for [62]). We also note that in contrast to other methods in the table, we are not using any supervision from a depth-from-single-image network.
Figure 5: $\gamma$ Visualization. We show a visualization of the $\gamma$ outputs of our network that are described in Sec. 4.2. In each video sequence, we show the input tracks, where each color visualizes its movement level value, $\gamma$. Purple marks rigid points with low $\gamma$ whereas red marks non-rigid points with high $\gamma$. Note, that our network did not get any direct supervision for these values, but the raw point tracks predictions from [18]. To emphasize our generalization, the $\gamma$ visualizations for cats were produced by the model that was only trained on dogs and vice versa.
5.4 Ablation Study
To evaluate the contribution of our different method parts we run an ablation study which is presented in Tab. 3. In this study, the training was always done on the cats partition from COP3D and evaluated on our test data which contains dogs and cats. More specifically, we ablate on our transformer architecture by taking the architecture suggested by [29] ("Set of Sets") or the DSS architecture [27] ("DSS"). As the table shows our architecture ("Full") achieved significantly better results. We further ignored the $\gamma$ outputs and used regular reprojection error on $B_{1}$ for all points ("No $\gamma$"). We further ablate our sparsity loss ("No $\mathcal{L}_{\text{Sparse}}$") and the rigid loss ("No $\mathcal{L}_{\text{Rigid}}$"). In both cases, the error increased whereas in the later one, the results became much worse.
We further ablate the choice of $K=12$ as the number of non-rigid bases, by trying two extreme numbers of $K=30,K=2$ (we saw no significant differences when we used nearby choices such as $K=11$). As can be seen in the table, when $K=30$ the output is not regularized enough and produces a higher depth error for the dynamic part. For $K=2$ the depth is regularized but the reprojection error ("Rep. (pix.)") gets higher due to over-regularization. Overall the ablation study justifies our design choices ("Full").
5.5 Limitations and Future Work
Our method cannot handle videos with too rapid motion, and in general, is limited by the accuracy of the tracking method [18]. We note that any future improvements with point tracking in terms of accuracy and inference time will directly improve our method as well.
Our method assumes enough motion parallax to constrain the depth values and fails to generate accurate camera poses without it. A future and interesting work would be to try combining depth-from-single-image prior as additional inputs to our network for handling cases with minimal motion parallax and for improving accuracies.
6 Conclusion
We presented TracksTo4D, a novel deep-learning framework that directly translates 2D motion tracks from casual videos into their corresponding non-rigid structure and motion. Our approach features a deep learning architecture that considers the symmetries in the problem with designed intrinsic constraints to handle the ill-posed nature of this problem. Notably, our network was trained using only raw supervision of 2D point tracks extracted by an off-the-shelf method [18] without any 3D supervision. Yet, it implicitly learned to predict camera poses and non-rigid structures while identifying the non-rigid parts. During inference, our method demonstrates faster processing times compared to previous methods while achieving comparable accuracy. Furthermore, our network exhibits strong generalization capabilities, performing well even on semantic categories that were not present in the training data.
Acknowledgments
HM is the Robert J. Shillman Fellow, and is supported by the Israel Science Foundation through a personal grant (ISF 264/23) and an equipment grant (ISF 532/23).
Appendix 0.A Supplementary Results
0.A.1 Video Results
We provide supplementary video outputs of several cases from our test set. Each video presents the input video frames with a set of pre-extracted point tracks that are used as input to our network and presented in corresponding colors (left side), and the output cameras and dynamic 3D structure (right side). The output camera trajectory is presented as gray frustums, whereas the current camera is marked in red. The reconstructed 3D scene points are presented in corresponding colors to the input tracks. Note that the outputs presented in the videos were obtained at inference time, with a single feed-forward prediction, without any optimization or fine-tuning, on unseen test cases.
0.A.2 Extended Quantitative Evaluation
Per-sequence and mean quantitative comparisons for our 21 pet test videos are presented in Tab. 5 and Tab. 6. We also show per-sequence fine-tuned results of 2000 iterations starting from our final checkpoint. The fine-tuning is done as a post-processing by minimizing the original loss function on the given test video. This takes about $10$ minutes per case which is still fast compared to the run time of the previous methods that predicted dynamic depths (see Tab. 1).
0.A.3 Generalization Beyond Pet Videos
We further applied our pretrained network on the MPI Sintel dataset [9] for the evaluation set defined by [64] of $14$ sequences. We note that the pet domain that was used for training our network is very different from the Sintel domain which contains fast motion of multiple characters. Yet, we can see that due to the strong prior collected by our network, it generalizes to some extent to the domain of Sintel. Fig. 6 shows a qualitative evaluation of our movement level predictions ($\gamma$). As can be seen, despite the domain gap (compared to Fig. 5), our network identifies the non-rigid movement in the videos. We further evaluated the camera pose accuracy in Tab. 7. Even though our mean accuracy is not the best among all methods, our method gets the most accurate results on some metrics for several sequences. Training on a video dataset in a similar domain to Sintel could possibly improve the generalization of our method for Sintel. Unfortunately, the Sintel dataset contains only 23 sequences.
Figure 6: $\gamma$ Visualization for Sintel. We show a visualization of the $\gamma$ outputs of our network. In each video sequence, we show the input tracks, where each color visualizes its movement level value, $\gamma$. Purple marks rigid points with low $\gamma$ whereas red marks non-rigid points with high $\gamma$. Note that our network was only trained on pet videos, and did not see any videos in the domain of Sintel. Yet, at inference time with no further optimization, it identifies the non-rigid movement successfully.
Appendix 0.B Implementation Details
Grid ATE $\downarrow$ RPE Trans $\downarrow$ RPE Rot $\downarrow$ Inference $\downarrow$
size (mm) (mm) (deg) Time
Ours (cats only) 15 8.96 3.79 0.23 0.16(+8.6) seconds
Ours (cats only)+BA 15 4.22 2.86 0.17 0.40(+8.6) seconds
Ours (cats only) 12 9.18 3.81 0.24 0.09(+7.8) seconds
Ours (cats only)+BA 12 4.36 2.97 0.17 0.24(+7.8) seconds
Ours (cats only) 10 8.91 3.91 0.23 0.05(+7.7) seconds
Ours (cats only)+BA 10 4.44 3.01 0.18 0.16(+7.7) seconds
Ours (cats only) 7 9.06 4.11 0.25 0.02(+7.6) seconds
Ours (cats only)+BA 7 4.93 3.52 0.20 0.08(+7.6) seconds
Ours (cats only) 5 10.29 4.97 0.31 0.01(+7.6) seconds
Ours (cats only)+BA 5 8.08 6.33 0.38 0.05(+7.6) seconds
Table 4: Tracking Grid Size Effect Quantitative evaluation of the effect of reducing the number of sampled point tracks at inference time. We measure the camera pose accuracy and the running time. We also mention the point tracks extraction time in parenthesis (e.g. +8.6 seconds) which is performed by [18] as a preprocess. As can be seen, our method can handle a smaller number of points but the accuracy slightly drops with fewer sampled points
0.B.1 Architecture
Here we provide a precise definition for our equivariant layer: $F:\mathbb{R}^{N\times P\times d}\rightarrow\mathbb{R}^{N\times P\times d}$. The computation of $F$ is split into four steps, as described below:
$\mathbf{\bar{q}}_{ij}=\bar{W}^{Q}M_{ij},~{}\mathbf{\bar{k}}_{ij}=\bar{W}^{K}M_{ij},~{}\mathbf{\bar{v}}_{ij}=\bar{W}^{V}M_{ij},$ (12a)
$\bar{M}_{ij}=\sum_{i^{\prime}=1}^{N}\frac{\exp(\mathbf{\bar{q}}_{ij}\cdot\mathbf{\bar{k}}_{i^{\prime}j})}{\sum_{l=1}^{N}\exp(\mathbf{\bar{q}}_{ij}\cdot\mathbf{\bar{k}}_{lj})}\mathbf{\bar{v}}_{i^{\prime}j}$ (12b)
$\mathbf{q}_{ij}=W^{Q}\bar{M}_{ij},~{}\mathbf{k}_{ij}=W^{K}\bar{M}_{ij},~{}\mathbf{v}_{ij}=W^{V}\bar{M}_{ij},$ (12c)
$F(M)_{ij}=\sum_{j^{\prime}=1}^{P}\frac{\exp(\mathbf{q}_{ij}\cdot\mathbf{k}_{ij^{\prime}})}{\sum_{l=1}^{P}\exp(\mathbf{q}_{ij}\cdot\mathbf{k}_{il})}\mathbf{v}_{ij^{\prime}}$ (12d)
Here, $M_{i,j}\in\mathbb{R}^{d}$ are the features associated with the $j$-th point in the $i$-th frame. The attention mechanism defined in the first two equations above (with additional temporal positional encoding) replaces the translation equivariant function $L$ applied to the columns of $M$ (Eq.(3)), and the attention in the last two equations above implements the point cloud aggregation (summation) (also in Eq.(3)) applied to the rows of $M$. As commonly done we use 16 attention head transformations [46]. We add standard temporal positional encoding for the transformers with respect to the $N$ (time) dimension as described in [46]. In our implementation, we added positional encoding only in the first layer.
Architecture Technical Details
For learning high frequencies we map each input coordinate to sinusoidal functions as in [28] with $L=12$. We use 3 pairs of attention layers, each of frames attention followed by point attention. Each point (after sinusoidal functions embedding) is mapped into $\mathbb{R}^{256}$ with a linear layer. Each attention layer is a function of the form $F:\mathbb{R}^{N\times P\times 256}\rightarrow\mathbb{R}^{N\times P\times 256}$ (see details above). Each attention uses 16 heads with $K,Q,V\in\mathbb{R}^{(N\text{ or }P)\times 64}$ followed by a fully connected network with 1 hidden layer of $2048$ features. We then average over the rows to get per-point features $P_{0}\in\mathbb{R}^{P\times 256}$ and over the columns to get per-frame features $F_{0}\in\mathbb{R}^{N\times 256}$. Finally, we map $P_{0}$ to per-point outputs $P_{1}\in\mathbb{R}^{P\times(3K+1)}$ ( K basis points and $\gamma$) with a linear layer, and $F_{0}$ into per-camera outputs $F_{1}\in\mathbb{R}^{N\times(6+3+K)}$ ($6$ for the rotation parameters [67], $3$ for the camera center, and K coefficients) using a conv layer with a kernel size of $31$.
0.B.2 Other Implementation Details
Point Tracks Sampling
For building $M\in\mathbb{R}^{N\times P\times 3}$ we use the implementation of [18]. We sample a uniform grid of $15\times 15$ 2D points, starting from frame number $0,20,40,\dots$, and then track these points throughout the entire video (backward and forward). In Tab. 4 we show the effect of reducing the grid size at inference time, in terms of camera pose accuracy and running time. During training, at each iteration, we randomly sample 20-50 frames from the training videos and 100 point tracks, i.e. $20\leq N\leq 50$ and $P=100$. When sampling cameras and point tracks of size $N\times P$ from a larger matrix of size $N^{\prime}\times P^{\prime}$ we only take a point track if its starting tracking time is in the range $[t-\frac{N}{2},t+\frac{3N}{2}]$, where t is the first sampled index. At inference time we take all the available point tracks. In both, training and inference time, we keep only point tracks that are observed in more than $10$ frames.
Test Set
We used the RGBD camera of the iPhone 11 to record our 21 test videos of dogs and cats. Each frame has a resolution of $640\times 480$ pixels. Note that the training set contained various types of resolutions. All pet owners who were photographed gave their permission for the animals to be photographed. For evaluation only, we define a point track as dynamic if its associated GT mask value is $1$ for at least $40$ frames. The GT masks are obtained by running [68] and searching for labels of dogs and cats. They were only used for evaluation and not used by our method at all.
Bundle Adjustment
After inference, as optional refinement, we take the output rigid approximation $B_{1}\in\mathbb{R}^{P\times 3}$ and the output camera poses $\{R_{1},\dots,R_{N}\}$, $\{\mathbf{t}_{1},\dots,\mathbf{t}_{N}\}$ and apply Bundle Adjustment (BA). We use a 3D world point from $B_{1}$ if its associated $\gamma$ is below $0.008$. We optimize reprojection errors of a given observation $M_{i,j}$, only if it is observed, i.e. $M_{i,j}^{o}=1$ and if the initial reprojection error is below $10$ pixels. We use the BA implementation provided by [29], which is based on the Ceres package [2].
Training Technical Details
In all training setups, we used: $\lambda_{\text{Reprojection}}=50.0$, $\lambda_{\text{Rigid}}=1.0$, $\lambda_{\text{Negative}}=1.0$, $\lambda_{\text{Sparse}}=0.001$. At the beginning of the training, we pre-train the camera poses to be located behind and facing the origin. This prevents cases in which the cameras are located in the middle of the initial point cloud s.t. many points have negative depths, which may result in bad convergence. More specifically, the pre-train loss is:
$\mathcal{L}_{\text{Pretrain}}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{100}\left\lVert\mathbf{t}_{i}-[0,0,-15]^{T}\right\rVert^{2}+\left\lVert R_{i}-I\right\rVert_{F}^{2}$
The pretrain runs until convergence ($\mathcal{L}_{\text{Pretrain}}<10^{-4})$. During the main training, we detach gradients from $B_{1}$ and $(R_{1},\mathbf{t}_{1})\dots,(R_{N},\mathbf{t}_{N})$ for $\mathcal{L}_{\text{Reproject}}$ to stabilize the training. Until epoch $50$ we sample sequences of $N$ in the range $[20,22]$, and then we increase the range to $[20,50]$.
Table 5: Depth Accuracy. We show a comparison to previous methods on the predicted depth for the point trajectories compared to their GT depths. We compare 4 ways of running our method. Ours (C&D), Ours (C): Our inference time outputs for a model that was trained on cats and dogs or only on cats respectively. Ours (C&D) FT, Ours (C) FT: The outputs of our model trained on cats and dogs or cats respectively, after fine-tuning our losses for each specific video. As can be seen, fine-tuning improves our accuracy even more.
RCVD [21] MiDaS[4] CasualSAM[62] Ours (C&D) Ours (C&D) FT Our (C) Our (C) FT
[…]
All 1.90E+08 8.80E+05 0.11 0.06 0.06 0.06 0.06
$\delta<1.25\uparrow$ Dyn 0.94 0.87 1.00 0.99 0.98 0.99 0.98
All 0.71 0.74 0.96 0.98 0.97 0.98 0.97
$\delta<1.25^{2}\uparrow$ Dyn 1.00 1.00 1.00 1.00 1.00 1.00 1.00
All 0.85 0.87 0.97 1.00 0.99 1.00 0.99
$\delta<1.25^{3}\uparrow$ Dyn 1.00 1.00 1.00 1.00 1.00 1.00 1.00
All 0.89 0.91 0.97 1.00 0.99 1.00 0.99
[…]
All 0.19 0.18 0.09 0.09 0.07 0.09 0.07
$\delta<1.25\uparrow$ Dyn 0.45 0.75 0.81 0.81 0.85 0.90 0.87
All 0.68 0.76 0.92 0.90 0.91 0.94 0.92
$\delta<1.25^{2}\uparrow$ Dyn 0.85 0.99 0.89 1.00 1.00 0.99 1.00
All 0.93 0.95 0.96 1.00 1.00 1.00 1.00
$\delta<1.25^{3}\uparrow$ Dyn 0.99 1.00 1.00 1.00 1.00 1.00 1.00
All 1.00 0.99 1.00 1.00 1.00 1.00 1.00
[…]
All 0.20 0.55 0.06 0.07 0.06 0.06 0.06
$\delta<1.25\uparrow$ Dyn 0.20 0.94 0.99 0.99 0.99 0.99 0.99
All 0.66 0.66 0.98 0.96 0.98 0.97 0.98
$\delta<1.25^{2}\uparrow$ Dyn 0.52 1.00 0.99 1.00 1.00 1.00 1.00
All 0.89 0.75 0.99 0.99 0.99 0.99 0.99
$\delta<1.25^{3}\uparrow$ Dyn 0.91 1.00 1.00 1.00 1.00 1.00 1.00
All 0.98 0.80 0.99 0.99 0.99 0.99 0.99
[…]
All 0.22 0.24 0.06 0.09 0.06 0.08 0.06
$\delta<1.25\uparrow$ Dyn 0.10 0.53 0.97 0.95 0.98 0.97 0.98
All 0.76 0.74 0.97 0.91 0.97 0.92 0.96
$\delta<1.25^{2}\uparrow$ Dyn 0.27 0.99 1.00 1.00 1.00 1.00 1.00
All 0.85 0.92 0.99 0.98 0.99 0.98 0.99
$\delta<1.25^{3}\uparrow$ Dyn 0.59 1.00 1.00 1.00 1.00 1.00 1.00
All 0.92 0.95 0.99 0.99 0.99 0.99 0.99
[…]
All 0.17 0.27 0.09 0.12 0.11 0.10 0.11
$\delta<1.25\uparrow$ Dyn 0.44 0.98 1.00 0.54 0.75 0.65 0.76
All 0.80 0.65 0.96 0.84 0.92 0.89 0.92
$\delta<1.25^{2}\uparrow$ Dyn 0.87 1.00 1.00 0.88 1.00 0.99 1.00
All 0.97 0.90 0.98 0.98 0.99 1.00 0.99
$\delta<1.25^{3}\uparrow$ Dyn 1.00 1.00 1.00 1.00 1.00 1.00 1.00
All 1.00 0.94 0.99 1.00 1.00 1.00 1.00
[…]
All 0.12 3.75E+05 0.03 0.08 0.04 0.06 0.04
$\delta<1.25\uparrow$ Dyn 0.97 0.98 1.00 0.99 0.98 0.99 0.99
All 0.86 0.86 0.98 0.91 0.98 0.96 0.98
$\delta<1.25^{2}\uparrow$ Dyn 1.00 1.00 1.00 1.00 1.00 1.00 1.00
All 0.97 0.95 1.00 0.97 1.00 0.98 1.00
$\delta<1.25^{3}\uparrow$ Dyn 1.00 1.00 1.00 1.00 1.00 1.00 1.00
All 0.99 0.98 1.00 0.99 1.00 1.00 1.00
[…]
All 0.14 0.15 0.05 0.04 0.06 0.05 0.06
$\delta<1.25\uparrow$ Dyn 0.47 0.94 0.98 0.99 0.99 0.99 0.99
All 0.83 0.87 0.96 0.98 0.96 0.97 0.96
$\delta<1.25^{2}\uparrow$ Dyn 0.65 0.99 0.99 1.00 1.00 1.00 1.00
All 0.91 0.95 0.99 0.99 0.98 0.99 0.98
$\delta<1.25^{3}\uparrow$ Dyn 0.99 0.99 1.00 1.00 1.00 1.00 1.00
All 0.99 0.98 1.00 1.00 0.99 1.00 0.99
[…]
All 0.22 0.17 0.06 0.10 0.06 0.09 0.06
$\delta<1.25\uparrow$ Dyn 0.45 0.95 0.92 0.84 0.89 0.90 0.91
All 0.65 0.80 0.97 0.86 0.96 0.91 0.96
$\delta<1.25^{2}\uparrow$ Dyn 0.68 1.00 0.99 0.98 0.99 0.98 0.99
All 0.86 0.92 0.99 0.98 0.99 0.98 0.99
$\delta<1.25^{3}\uparrow$ Dyn 0.93 1.00 1.00 1.00 1.00 1.00 1.00
All 0.97 0.98 1.00 0.99 1.00 0.99 1.00
[…]
All 0.20 3.80E+04 0.03 0.08 0.04 0.07 0.04
$\delta<1.25\uparrow$ Dyn 0.18 0.68 0.99 0.97 0.99 0.98 0.99
All 0.72 0.64 0.99 0.89 0.99 0.97 0.99
$\delta<1.25^{2}\uparrow$ Dyn 0.69 0.97 1.00 1.00 1.00 1.00 1.00
All 0.91 0.84 1.00 0.99 1.00 1.00 1.00
$\delta<1.25^{3}\uparrow$ Dyn 0.97 1.00 1.00 1.00 1.00 1.00 1.00
All 0.99 0.92 1.00 1.00 1.00 1.00 1.00
[…]
All 0.33 0.20 0.09 0.21 0.09 0.22 0.09
$\delta<1.25\uparrow$ Dyn 0.05 0.59 0.80 0.46 0.60 0.45 0.59
All 0.65 0.74 0.92 0.55 0.88 0.52 0.88
$\delta<1.25^{2}\uparrow$ Dyn 0.28 0.85 0.95 0.91 1.00 0.85 1.00
All 0.78 0.91 0.98 0.89 0.99 0.86 0.99
$\delta<1.25^{3}\uparrow$ Dyn 0.68 1.00 1.00 1.00 1.00 1.00 1.00
All 0.90 0.98 1.00 0.97 1.00 0.97 1.00
[…]
All 0.11 0.24 0.03 0.03 0.02 0.04 0.02
$\delta<1.25\uparrow$ Dyn 0.99 1.00 1.00 1.00 1.00 1.00 1.00
All 0.92 0.73 1.00 1.00 1.00 0.99 1.00
$\delta<1.25^{2}\uparrow$ Dyn 1.00 1.00 1.00 1.00 1.00 1.00 1.00
All 0.99 0.87 1.00 1.00 1.00 1.00 1.00
$\delta<1.25^{3}\uparrow$ Dyn 1.00 1.00 1.00 1.00 1.00 1.00 1.00
All 1.00 0.95 1.00 1.00 1.00 1.00 1.00
[…]
All 0.17 0.38 0.05 0.08 0.05 0.07 0.05
$\delta<1.25\uparrow$ Dyn 0.47 0.79 0.92 0.90 0.92 0.90 0.92
All 0.74 0.67 0.97 0.90 0.96 0.94 0.96
$\delta<1.25^{2}\uparrow$ Dyn 0.86 1.00 0.99 0.96 0.98 0.97 0.98
All 0.94 0.78 0.99 0.99 1.00 0.99 0.99
$\delta<1.25^{3}\uparrow$ Dyn 0.94 1.00 1.00 0.99 1.00 0.99 1.00
All 0.99 0.86 1.00 1.00 1.00 1.00 1.00
[…]
All 0.16 0.13 0.05 0.08 0.08 0.08 0.08
$\delta<1.25\uparrow$ Dyn 0.19 0.79 0.98 0.98 0.97 0.98 0.97
All 0.73 0.85 0.98 0.95 0.94 0.95 0.94
$\delta<1.25^{2}\uparrow$ Dyn 0.97 1.00 1.00 1.00 1.00 1.00 1.00
All 0.98 0.98 1.00 0.99 0.97 0.99 0.97
$\delta<1.25^{3}\uparrow$ Dyn 1.00 1.00 1.00 1.00 1.00 1.00 1.00
All 1.00 1.00 1.00 1.00 0.98 1.00 0.98
[…]
All 3.31E+08 0.35 0.06 0.10 0.08 0.09 0.07
$\delta<1.25\uparrow$ Dyn 0.50 0.50 0.93 0.82 0.95 0.85 0.96
All 0.67 0.69 0.97 0.91 0.97 0.93 0.97
$\delta<1.25^{2}\uparrow$ Dyn 0.81 0.84 1.00 1.00 1.00 1.00 1.00
All 0.81 0.87 0.99 0.98 0.99 0.99 0.99
$\delta<1.25^{3}\uparrow$ Dyn 1.00 1.00 1.00 1.00 1.00 1.00 1.00
All 0.87 0.93 1.00 0.99 0.99 0.99 0.99
[…]
All 0.18 0.34 0.03 0.04 0.03 0.03 0.03
$\delta<1.25\uparrow$ Dyn 0.79 0.72 0.99 0.99 0.99 0.99 0.99
All 0.76 0.55 1.00 0.99 0.99 0.99 0.99
$\delta<1.25^{2}\uparrow$ Dyn 0.93 0.97 1.00 1.00 1.00 1.00 1.00
All 0.94 0.83 1.00 1.00 1.00 1.00 1.00
$\delta<1.25^{3}\uparrow$ Dyn 0.99 1.00 1.00 1.00 1.00 1.00 1.00
All 0.99 0.92 1.00 1.00 1.00 1.00 1.00
[…]
All 0.33 0.39 0.09 0.15 0.11 0.18 0.10
$\delta<1.25\uparrow$ Dyn 0.01 0.77 0.95 0.62 0.90 0.79 0.88
All 0.65 0.65 0.97 0.80 0.93 0.70 0.93
$\delta<1.25^{2}\uparrow$ Dyn 0.12 1.00 0.99 0.93 0.99 0.96 0.99
All 0.81 0.83 0.99 0.97 0.98 0.92 0.98
$\delta<1.25^{3}\uparrow$ Dyn 0.46 1.00 0.99 1.00 0.99 1.00 0.99
All 0.90 0.89 0.99 0.99 0.98 0.99 0.99
[…]
All 0.21 0.21 0.05 0.06 0.06 0.06 0.06
$\delta<1.25\uparrow$ Dyn 0.42 0.92 0.98 0.81 0.89 0.80 0.90
All 0.66 0.79 0.99 0.96 0.97 0.96 0.97
$\delta<1.25^{2}\uparrow$ Dyn 0.98 1.00 1.00 1.00 1.00 1.00 1.00
All 0.95 0.93 1.00 1.00 1.00 1.00 1.00
$\delta<1.25^{3}\uparrow$ Dyn 1.00 1.00 1.00 1.00 1.00 1.00 1.00
All 1.00 0.96 1.00 1.00 1.00 1.00 1.00
[…]
All 0.23 0.15 0.08 0.07 0.07 0.07 0.06
$\delta<1.25\uparrow$ Dyn 0.40 0.88 0.76 0.93 0.87 0.91 0.88
All 0.60 0.81 0.89 0.94 0.94 0.93 0.94
$\delta<1.25^{2}\uparrow$ Dyn 0.82 0.99 0.98 0.99 0.98 0.98 0.98
All 0.87 0.94 0.98 0.98 0.99 0.99 0.99
$\delta<1.25^{3}\uparrow$ Dyn 0.95 1.00 1.00 0.99 1.00 0.99 1.00
All 0.95 0.98 1.00 0.99 1.00 1.00 1.00
[…]
All 0.14 0.32 0.05 0.10 0.09 0.08 0.09
$\delta<1.25\uparrow$ Dyn 0.46 0.60 0.93 0.87 0.95 0.88 0.95
All 0.86 0.65 0.98 0.91 0.96 0.95 0.96
$\delta<1.25^{2}\uparrow$ Dyn 0.55 0.95 0.98 1.00 0.99 0.99 0.99
All 0.91 0.84 0.99 0.98 0.98 0.98 0.98
$\delta<1.25^{3}\uparrow$ Dyn 0.84 1.00 0.99 1.00 1.00 1.00 1.00
All 0.97 0.92 1.00 0.99 0.98 0.99 0.98
[…]
All 0.15 0.27 0.03 0.03 0.03 0.05 0.03
$\delta<1.25\uparrow$ Dyn 0.28 0.80 0.97 0.97 0.97 0.98 0.97
All 0.81 0.74 0.99 0.99 0.99 0.99 0.99
$\delta<1.25^{2}\uparrow$ Dyn 0.91 0.97 1.00 1.00 0.99 0.99 0.99
All 0.98 0.87 1.00 1.00 1.00 1.00 1.00
$\delta<1.25^{3}\uparrow$ Dyn 1.00 0.99 1.00 1.00 1.00 1.00 1.00
All 1.00 0.92 1.00 1.00 1.00 1.00 1.00
[…]
All 2.42E+08 0.45 0.04 0.05 0.05 0.05 0.05
$\delta<1.25\uparrow$ Dyn 0.34 0.49 0.61 0.44 0.38 0.49 0.34
All 0.50 0.42 0.97 0.95 0.95 0.95 0.95
$\delta<1.25^{2}\uparrow$ Dyn 0.93 0.90 1.00 1.00 1.00 1.00 1.00
All 0.77 0.72 1.00 1.00 1.00 1.00 1.00
$\delta<1.25^{3}\uparrow$ Dyn 0.99 0.99 1.00 1.00 1.00 1.00 1.00
All 0.83 0.85 1.00 1.00 1.00 1.00 1.00
[…]
All 3.63E+07 6.16E+04 0.06 0.08 0.06 0.08 0.06
$\delta<1.25\uparrow$ Dyn 0.43 0.78 0.93 0.85 0.90 0.88 0.89
All 0.72 0.71 0.97 0.91 0.96 0.92 0.96
$\delta<1.25^{2}\uparrow$ Dyn 0.75 0.97 0.99 0.98 1.00 0.99 1.00
All 0.90 0.88 0.99 0.98 0.99 0.98 0.99
$\delta<1.25^{3}\uparrow$ Dyn 0.92 1.00 1.00 1.00 1.00 1.00 1.00
All 0.96 0.93 1.00 1.00 0.99 1.00 0.99
Table 6: Camera Poses Accuracy. We show a comparison to previous methods on the predicted camera poses. We compare 3 ways of running our method. Ours (C): Our inference time outputs (total inference time of 0.16 seconds) for a model that was trained only on cats. Ours (C)+BA: Our inference time outputs, followed by a short Bundle Adjustment (total inference time of 0.4 seconds) for a model that was trained only on cats. Ours (C) FT: The outputs of the model that was trained only on cats after fine-tuning our losses for each specific video (total running time of about 10 minutes). As can be seen, after BA, our results are the most accurate compared to the other methods, and fine-tuning improves our accuracy even more.
DROID-SLAM[43] ParticleSfM[64] RCVD[21] CasualSAM[62] Ours Ours Ours
(C) (C)+BA (Csho)+FT
Seq0 ATE(mm) 3.71 6.10 64.67 5.36 5.60 5.43 4.40
RPE T.(mm) 3.05 3.22 26.92 3.13 3.63 3.04 2.91
RPE R.(deg.) 0.14 0.18 2.53 0.16 0.22 0.15 0.13
Seq1 ATE(mm) 1.91 3.83 38.44 10.28 13.21 4.32 1.71
RPE T.(mm) 1.32 1.49 25.23 3.00 2.65 1.72 1.29
RPE R.(deg.) 0.18 0.23 2.43 0.67 0.29 0.19 0.16
Seq2 ATE(mm) 3.13 4.68 39.46 2.13 4.57 2.78 2.52
RPE T.(mm) 3.62 5.82 27.27 1.97 3.24 2.62 2.55
RPE R.(deg.) 0.08 0.21 1.89 0.06 0.12 0.09 0.09
Seq3 ATE(mm) 5.13 2.26 29.55 2.49 7.83 2.66 2.87
RPE T.(mm) 5.76 2.00 24.18 2.31 3.13 2.27 2.38
RPE R.(deg.) 0.17 0.07 2.38 0.08 0.13 0.07 0.07
Seq4 ATE(mm) 2.59 4.38 56.16 2.65 4.21 2.38 2.24
RPE T.(mm) 2.05 2.07 21.36 1.74 2.40 2.05 1.95
RPE R.(deg.) 0.11 0.11 1.05 0.09 0.15 0.11 0.10
Seq5 ATE(mm) 1.07 0.83 14.22 0.53 1.44 0.79 0.74
RPE T.(mm) 1.02 0.74 6.51 0.52 0.88 0.68 0.68
RPE R.(deg.) 0.09 0.07 1.63 0.05 0.12 0.07 0.06
Seq6 ATE(mm) 26.08 31.07 48.31 26.12 27.10 28.28 29.80
RPE T.(mm) 10.57 11.01 24.21 10.31 10.44 10.92 11.02
RPE R.(deg.) 0.66 0.67 4.03 0.62 0.69 0.69 0.68
Seq7 ATE(mm) 2.25 28.48 47.25 1.78 4.53 2.39 2.11
RPE T.(mm) 2.34 6.13 23.81 1.72 2.32 1.98 1.95
RPE R.(deg.) 0.15 0.59 2.29 0.11 0.20 0.16 0.15
Seq8 ATE(mm) 1.23 38.79 44.06 2.00 4.78 0.84 0.96
RPE T.(mm) 1.07 50.57 25.46 1.24 1.54 0.71 0.74
RPE R.(deg.) 0.07 4.78 1.73 0.09 0.13 0.06 0.05
Seq9 ATE(mm) 21.74 18.52 43.45 34.93 24.15 3.92 3.28
RPE T.(mm) 9.04 7.96 21.35 21.06 6.89 2.77 2.01
RPE R.(deg.) 0.62 0.47 3.47 0.71 0.37 0.12 0.10
Seq10 ATE(mm) 1.47 1.75 22.49 1.40 3.03 1.42 1.48
RPE T.(mm) 1.84 1.15 24.22 1.11 1.87 1.25 1.24
RPE R.(deg.) 0.22 0.12 2.18 0.10 0.21 0.12 0.12
Seq11 ATE(mm) 1.71 3.60 19.10 1.32 3.12 1.66 1.50
RPE T.(mm) 1.65 1.54 16.34 1.21 3.12 1.75 1.60
RPE R.(deg.) 0.08 0.08 1.80 0.06 0.12 0.08 0.07
Seq12 ATE(mm) 1.70 3.64 20.82 2.51 6.23 3.61 2.58
RPE T.(mm) 2.24 2.63 18.50 2.74 4.04 3.02 2.34
RPE R.(deg.) 0.09 0.12 2.03 0.12 0.23 0.15 0.14
Seq13 ATE(mm) 1.23 2.38 33.49 1.49 4.10 1.17 1.27
RPE T.(mm) 1.19 1.40 17.33 1.00 1.72 1.12 1.01
RPE R.(deg.) 0.13 0.14 2.12 0.12 0.19 0.12 0.11
Seq14 ATE(mm) 5.42 5.15 1.05E+02 24.95 6.09 5.06 4.90
RPE T.(mm) 3.40 3.38 69.36 9.45 4.07 3.57 3.55
RPE R.(deg.) 0.18 0.19 3.81 0.65 0.26 0.20 0.19
Seq15 ATE(mm) 7.69 61.06 36.70 7.40 36.51 4.98 5.50
RPE T.(mm) 7.95 17.57 28.18 6.93 10.19 5.86 5.74
RPE R.(deg.) 0.22 0.58 2.85 0.19 0.30 0.16 0.15
Seq16 ATE(mm) 5.04 5.06 36.42 3.69 4.52 4.11 3.81
RPE T.(mm) 4.53 4.54 20.11 4.02 4.47 4.25 4.25
RPE R.(deg.) 0.28 0.29 2.94 0.25 0.31 0.28 0.28
Seq17 ATE(mm) 1.12 34.07 77.91 2.08 6.52 2.51 2.89
RPE T.(mm) 1.15 12.30 39.64 1.18 2.14 1.50 1.47
RPE R.(deg.) 0.10 1.11 2.20 0.08 0.17 0.13 0.13
Seq18 ATE(mm) 2.98 6.25 36.54 4.91 7.73 4.13 3.83
RPE T.(mm) 4.05 5.21 24.78 3.89 4.34 3.62 3.80
RPE R.(deg.) 0.23 0.29 1.67 0.21 0.24 0.21 0.21
Seq19 ATE(mm) 1.45 2.23 42.95 1.82 8.51 2.79 2.39
RPE T.(mm) 1.65 1.72 29.12 1.44 3.23 2.21 1.97
RPE R.(deg.) 0.11 0.11 2.09 0.09 0.25 0.16 0.14
Seq20 ATE(mm) 8.13 4.37 66.03 4.95 4.30 3.43 4.03
RPE T.(mm) 6.20 3.57 27.34 3.00 3.30 3.05 3.08
RPE R.(deg.) 0.36 0.20 1.35 0.18 0.20 0.18 0.18
Mean ATE(mm) 5.08 12.79 43.95 6.90 8.96 4.22 4.04
RPE T.(mm) 3.60 6.95 25.77 3.95 3.79 2.86 2.74
RPE R.(deg.) 0.20 0.51 2.31 0.22 0.23 0.17 0.16
Table 7: Camera Poses Accuracy on Sintel. We show a comparison to previous methods on the predicted camera poses. We compare 4 ways of running our method. Ours (C), Ours (C&D): Our inference time outputs for a model that was trained only on cats or on cats and dogs respectively. Ours (C)+BA, Ours (C&D)+BA: The outputs of the model trained only on cats or on cats and dogs respectively, after further applying short BA refinement. The numbers for [21],[43] and [64] were obtained from [64].
RCVD[21] DROID- Particle CasualSAM[62] Ours Ours Ours Ours
SLAM[43] SfM[64] (C) (C)+BA (C&D) (C&D)+BA
alley_2 ATE(mm) 0.026 0.057 0.002 0.0008 0.043 0.027 0.055 0.046
RPE T.(mm) 0.056 0.035 0.0007 0.0004 0.033 0.040 0.034 0.041
RPE R.(deg.) 0.821 1.047 0.009 0.008 0.468 0.507 0.511 0.531
ambush_4 ATE(mm) 0.171 0.104 0.068 0.011 0.050 0.021 0.055 0.022
RPE T.(mm) 0.048 0.035 0.027 0.007 0.026 0.026 0.026 0.026
RPE R.(deg.) 3.025 1.385 0.473 0.079 0.398 0.319 0.382 0.324
ambush_5 ATE(mm) 0.230 0.112 0.002 0.008 0.057 0.010 0.035 0.009
RPE T.(mm) 0.046 0.029 0.001 0.003 0.018 0.008 0.015 0.006
RPE R.(deg.) 4.105 1.580 0.055 0.084 0.430 0.228 0.338 0.149
ambush_6 ATE(mm) 0.199 0.289 0.269 0.042 0.241 0.223 0.191 0.141
RPE T.(mm) 0.112 0.078 0.081 0.024 0.088 0.129 0.072 0.078
RPE R.(deg.) 4.147 4.596 0.951 0.482 1.928 3.364 1.602 1.849
cave_2 ATE(mm) 0.596 0.351 0.961 0.967 0.217 0.217 0.460 0.460
RPE T.(mm) 0.171 0.172 0.142 0.170 0.094 0.094 0.113 0.113
RPE R.(deg.) 7.508 5.489 3.678 2.674 1.781 1.781 2.634 2.634
cave_4 ATE(mm) 0.179 0.155 0.068 0.009 0.055 0.006 0.115 0.025
RPE T.(mm) 0.087 0.035 0.012 0.003 0.024 0.006 0.028 0.021
RPE R.(deg.) 2.040 2.710 0.409 0.052 0.244 0.044 0.294 0.171
market_2 ATE(mm) 0.032 0.011 0.003 0.001 0.027 0.002 0.033 0.003
RPE T.(mm) 0.018 0.006 0.010 0.0006 0.004 0.002 0.007 0.002
RPE R.(deg.) 0.141 0.036 0.041 0.010 0.107 0.008 0.091 0.008
market_5 ATE(mm) 1.213 0.912 0.012 0.005 0.891 1.187 0.859 0.525
RPE T.(mm) 0.762 0.293 0.006 0.002 0.212 0.557 0.209 0.395
RPE R.(deg.) 1.863 3.334 0.034 0.016 1.252 6.946 1.489 3.285
market_6 ATE(mm) 0.248 0.057 0.018 0.008 0.187 0.163 0.220 0.152
RPE T.(mm) 0.214 0.037 0.007 0.002 0.103 0.097 0.095 0.086
RPE R.(deg.) 0.817 1.296 0.120 0.019 1.003 0.850 0.834 0.695
shaman_3 ATE(mm) 0.054 0.001 0.0005 0.002 0.002 0.002 0.003 0.002
RPE T.(mm) 0.023 0.002 0.0004 0.0004 0.0007 0.0004 0.001 0.0004
RPE R.(deg.) 0.718 0.199 0.072 0.036 0.060 0.042 0.087 0.039
sleeping_1 ATE(mm) 0.029 0.011 0.008 0.008 0.002 0.002 0.002 0.002
RPE T.(mm) 0.019 0.006 0.001 0.001 0.001 0.0006 0.001 0.0006
RPE R.(deg.) 0.668 0.479 0.042 0.040 0.070 0.040 0.096 0.040
sleeping_2 ATE(mm) 0.043 0.005 0.0007 0.0002 0.005 0.0006 0.005 0.0006
RPE T.(mm) 0.049 0.018 0.0002 0.0002 0.003 0.0007 0.003 0.0007
RPE R.(deg.) 0.446 0.139 0.006 0.008 0.108 0.018 0.088 0.018
temple_2 ATE(mm) 1.245 0.073 0.011 0.086 0.189 0.011 0.217 0.009
RPE T.(mm) 0.394 0.348 0.002 0.008 0.029 0.004 0.032 0.003
RPE R.(deg.) 1.318 1.298 0.019 0.131 0.287 0.019 0.340 0.018
temple_3 ATE(mm) 0.769 0.310 0.381 0.587 0.350 0.510 0.329 0.473
RPE T.(mm) 0.161 0.093 0.149 0.136 0.114 0.190 0.114 0.138
RPE R.(deg.) 20.592 3.230 1.583 0.109 1.672 2.110 1.458 1.176
Mean ATE(mm) 0.360 0.175 0.129 0.124 0.165 0.170 0.184 0.133
RPE T.(mm) 0.154 0.085 0.031 0.025 0.054 0.082 0.054 0.065
RPE R.(deg.) 3.444 1.916 0.535 0.268 0.701 1.163 0.732 0.781

### v2

@@
-Learning Priors for Non Rigid SfM from Casual Videos
+Fast Encoder-Based 3D from Casual Videos via Point Track Processing
 Abstract
-We tackle the long-standing challenge of reconstructing 3D structures and camera positions from videos.
-The problem is particularly hard when objects are transformed in a non-rigid way.
-Current approaches to this problem make unrealistic assumptions or require a long optimization time.
-We present TracksTo4D, a novel deep learning-based approach that enables inferring 3D structure and camera positions from dynamic content originating from in-the-wild videos using a single feed-forward pass on a sparse point track matrix.
-To achieve this, we leverage recent advances in 2D point tracking and design an equivariant neural architecture tailored for directly processing 2D point tracks by leveraging their symmetries.
-TracksTo4D is trained on a dataset of in-the-wild videos utilizing only the 2D point tracks extracted from the videos, without any 3D supervision.
-Our experiments demonstrate that TracksTo4D generalizes well to unseen videos of unseen semantic categories at inference time, producing equivalent results to state-of-the-art methods while significantly reducing the runtime compared to other baselines.
+This paper addresses the long-standing challenge of reconstructing 3D structures from videos with dynamic content.
+Current approaches to this problem were not designed to operate on casual videos recorded by standard cameras or require a long optimization time.
+Aiming to significantly improve the efficiency of previous approaches, we present TracksTo4D, a learning-based approach that enables inferring 3D structure and camera positions from dynamic content originating from casual videos using a single efficient feed-forward pass.
+To achieve this, we propose operating directly over 2D point tracks as input and designing an architecture tailored for processing 2D point tracks.
+Our proposed architecture is designed with two key principles in mind: (1) it takes into account the inherent symmetries present in the input point tracks data, and (2) it assumes that the movement patterns can be effectively represented using a low-rank approximation.TracksTo4D is trained in an unsupervised way on a dataset of casual videos utilizing only the 2D point tracks extracted from the videos, without any 3D supervision.
+Our experiments show that TracksTo4D can reconstruct a temporal point cloud and camera positions of the underlying video with accuracy comparable to state-of-the-art methods, while drastically reducing runtime by up to 95%.
+We further show that TracksTo4D  generalizes well to unseen videos of unseen semantic categories at inference time.
 Project page: https://tracks-to-4d.github.io.
-Keywords:
-Structure From Motion Dynamic Videos 3D Reconstruction Equivariance Symmetries
+1 Introduction
+Predicting 3D geometry in dynamic scenes is a challenging problem.
+In this problem setup, we are given access to multiple images of a scene taken sequentially, e.g., from a monocular video camera, where both the content in the scene and the camera are moving.
+Our task is to reconstruct the dynamic 3D positions of the points seen in the images and the camera poses.
+This fundamental problem has gained significant interest from the research community over the years [6, 34, 23, 59], mainly due to its important applications in many fields such as robot navigation, autonomous driving and 3D reconstruction of general environments [17].
+Importantly, in contrast to static scenes where the epipolar geometry constraints hold between the corresponding points of different views [15], determining the depth of a moving point from monocular views is an ill-posed problem [3].
+This causes standard Structure from Motion techniques [38, 51, 30] to be inadequate in this setup [22].
+Previous work and limitations.
+Many existing approaches for the above problem make simplifying assumptions that limit their applicability to real-world scenarios.
+Methods based on orthographic camera models and low-rank assumptions use matrix factorization techniques [6, 23], but the orthographic camera assumption might not be realistic and may cause reconstruction errors.
+Techniques that incorporate depth priors often require lengthy optimization processes in order to make the depth estimates across frames consistent [22, 59].
+Other physics-based approaches make assumptions about rigid bones [56, 54] or isometric deformable surfaces [34] and typically involve complex, slow optimization per video.
+In addition, they may require foreground-background segmentation of the moving content, which is not always easily obtained.
+Alternatively, some methods are specifically tailored to certain object classes like humans [48], restricting their domain to those limited cases.
+Consequently, these prior methods are either not directly applicable to casual videos, or require long optimization time per video.
 Figure 1: We present TracksTo4D, a method for mapping a set of 2D point tracks extracted from casual dynamic videos into their corresponding 3D locations and camera motion.
-At inference time, our network predicts the non-rigid structure and camera motion in a single feed-forward pass.
-Top.
-An example showing the input video frames ordered from left to right.
-A set of pre-extracted point tracks that are used as input to our network and presented in corresponding colors.
-Bottom.
-The outputs of our network for 3 time steps as observed from the right side of the scene.
-The camera trajectory is present as gray frustums, whereas the current camera is marked in red.
-The reconstructed 3D scene points are presented in corresponding colors to the input tracks on the top.
-See video visualizations here.
-1 Introduction
-Predicting 3D structures in dynamic scenes is a challenging problem.
-In contrast to rigid scenes where the epipolar geometry constraints hold between the corresponding points of different views [17], determining the depth of a moving point from monocular views is an ill-posed problem.
-Moreover, while multiview camera pose estimation can be solved successfully by Structure from Motion methods [37, 53, 30], these techniques tend to fail in scenarios with significant non-rigid motion [21].
-Background and previous work.
-Unfortunately, current approaches to solving this problem are based on unrealistic assumptions or suffer from long optimization times per scene that make them inapplicable to real-world problems.
-More concretely, early techniques for Non-Rigid SFM [6] used matrix factorization of long-term point tracks matrix of the non-rigid scene part.
-This was done by assuming that a small linear basis and per-frame coefficients can formulate the dynamic structure.
-By further assuming an orthographic camera is used, they omitted camera translations and solved for rotations and 3D structure by matrix factorization.
-While follow-up work improved accuracies [22, 11], this setup is still not practical for real-world casual videos, captured by pinhole cameras.
-Recently, the Neural Radiance Field (NeRF) [28] concept has brought major progress to novel view synthesis and surface reconstruction [60, 49] for static scenes.
-While some follow-up work extended NeRF to non-rigid scenes [23, 24, 33], optimizing a non-rigid NeRF requires a long, and per-scene optimization.
-Other recent approaches [26, 21, 62, 63] optimized for consistent depth maps given a specific monocular video by fine-tuning pre-trained single-view-depth-predictions [34].
-These methods showed impressive progress in camera and non-rigid structure estimation, but also require per-video optimization at inference time.
-Other approaches learned prior and parameterizations for specific non-rigid object classes such as humans [36, 25].
-This prior enables shorter inference times for obtaining 3D structures from monocular videos that contain instances from the learned category [55].
-However, this approach does not generalize well across shape categories and mostly requires per-video finetuning.
-Our approach.
-In this work, we propose TracksTo4D, a novel deep learning-based approach for predicting the 3D dynamic Non-Rigid Structure And camera Motion from videos.
-Our key observation is that the 2D motion present in casual videos shares common patterns across a variety of semantic video categories.
-Crucially, and as was suggested by [6], this motion can be effectively encoded into a point track matrix, where each entry represents the 2D position of a tracked point in a specific video frame.
-Rather than processing image pixels or semantic features that vary dramatically across different scenes and objects, we propose directly processing these point track matrices to learn generic, class-agnostic features.
-This also differs from the common approach of using optical flow representations of motion, as utilizing optical flow fields as input to deep neural networks is non-trivial.
-TracksTo4D  builds on the recent progress in point tracking [12, 18], which introduced the ability to extract long-term point tracks from casual videos.
-We leverage this progress and design a deep neural network architecture, that maps the point tracks of a given video into corresponding camera poses, 3D points, and per-point non-rigid motion level values (See Fig.
-1).
-Crucially, the mapping we aim to learn preserves two natural symmetries: (1) the points being tracked can be arbitrarily permuted without affecting the problem; (2) the frames containing these points exhibit temporal structure, adhering to an approximate time-translation symmetry.
-Following the Geometric Deep Learning paradigm [7], we build upon recent theoretical advances in equivariant learning [27] and integrate these two symmetries into our network architecture using dedicated attention mechanisms.
-Predicting the camera poses and the 3D non-rigid motion from 2D motion alone is an ill-posed problem without any constraints, and equivalent to finding the 3D structure from each frame [3].
-Inspired by previous work [6], we model the predicted 3D point cloud as a weighted sum of a 3D structure basis, where the first basis is assumed to fully represent the rigid points, and the rest of the basis elements are used to complete the non-rigid deviation for the non-rigid points.
-The predicted motion level values determine the non-rigidity of each 3D world point, i.e.
-the inconsistency of the rigid assumption that is represented by the first basis element for this world point.
-In our experiments, TracksTo4D is trained on the Common Pets dataset [38] using only 2D point tracks extracted by [18], without any 3D supervision by simply minimizing the reprojection errors.
-Interestingly, we observe that by just "watching" point tracks in videos, our network learns to infer their 3D location and the camera motion.
-We evaluate our method on test data with GT depth maps and demonstrate that it generalizes well, and produces comparable results to the state-of-the-art methods, with significantly shorter inference time.
-Contributions.
-In summary, our contributions are: (1) Designing an equivariant network architecture that maps from 2D point trajectories into corresponding non-rigid structures and camera poses; (2) Introducing a training paradigm to identify the rigid part of the scene for constraining the camera poses; (3) Simple training on casual videos without any 3D supervision; (4) Demonstrating fast inference time, accurate results, and strong generalization across animal categories.
-2 Related Work
-Structure from Motion
-The Structure from Motion (SfM) problem recovers the 3D structure and the camera poses from images.
-Early methods [44, 39] factorized a 2D measurement matrix of corresponding image points into a multiplication of a camera poses matrix by a 3D points matrix.
-These methods assume that each recovered 3D point is tracked over the entire image sequence which is not the case in practical use cases.
-Incremental approaches for SfM [37, 1, 53] start from a minimal set of images, to build an initial reconstruction.
-Then, the rest of the images are added to the scene one by one, by finding 2D correspondences from the new image into scene points.
-To prevent drift, Bundle Adjustment [45] is repeatedly applied to the reconstruction during this process.
-Several global methods [52, 10, 19, 51, 32] reduced the running time by using relative two-view geometries between image pairs to extract absolute camera poses in a single step, followed by a final round of Bundle Adjustment.
-Recently, Moran et al.
-[29] presented a method for applying an order equivariant architecture for mapping a set of multiview image correspondences into two sets of cameras and 3D points.
-It was trained on a dataset of scenes and obtained a reduced optimization time on novel scenes, due to the learned prior.
-A follow-up work [8] improved generalization by adding graph attention layers into the architecture of Moran et al.
-Other recent deep learning-based techniques used differentiable Bundle Adjustment layers [47, 40], or diffusion models [61, 48].
-The above-mentioned techniques assumed that the structure is rigid such that the whole scene can be represented in a single 3D structure.
-Simultaneous Localization and Mapping (SLAM)
-In contrast to SfM which assumes unordered images, monocular Simultaneous Localization and Mapping (SLAM) methods [30, 31, 15] get as input a sequence of video frames recorded by a single camera.
-These methods often define a set of keyframes for reconstructing the scene with a mapping thread, while registering every new frame into the structure by a localization thread.
-The mapping thread repeatedly applies Bundle Adjustment for updating the structure.
-ORB-SLAM [30] uses feature points for mapping a sparse scene representation and tracks the camera with sparse point matches.
-Other methods [31, 15, 14] use denser scene representations and align cameras directly by using photometric constraints.
-More recent methods combine classical approaches with deep learning-based tools e.g.
-depth prediction from a single image, optical flow prediction, or relative camera pose estimation [59, 5, 50, 65, 66, 41].
-As in SfM, these methods assume rigid scenes, fail to produce the cameras in scenes with large portions of non-rigid motion, and cannot reproduce non-rigid parts of the scene.
-DROID-SLAM [43] used synthetic data with ground truth 3D supervision for learning to predict camera poses via deep-based bundle adjustment on keyframes while excluding dynamic objects.
-ParticleSfM [64] pre-trained a motion prediction network on a synthetic dataset with ground truth supervision, and used it to filter out any non-rigid content for reproducing the cameras in dynamic scenes.
-Both, [43, 64] do not infer the non-rigid 3D structure.
-Test-Time Optimization for Non-Rigid Scenes
-Recent methods [26, 21, 62, 63] finetuned the monocular depth estimation from a pre-trained model [35, 34] using optical flow constraints [42], for obtaining consistent dense depth maps for a monocular video.
-[62] further optimized motion maps that represent the $\gamma$ parameters of zero mean Cauchy distributions for handling scenes with large dynamic motion.
-[54, 16] use depth from single image estimations, to improve novel view synthesis in dynamic scenes.
-[24] further optimized for the unknown camera poses together with the dynamic radiance field optimization.
-LASR [56], ViSER [57] and BANMo [58] optimize for a dynamic surface given a monocular video by assuming rigid bones and linear blend skinning weights.
-However, all the above-mentioned methods require per-scene optimization, resulting in slow inference.
-Recently, [38] presented the Common Pets in 3D (COP3D) dataset that contains casual, in-the-wild videos of pets, and used it to learn priors for novel view synthesis in dynamic scenes.
-3 Problem formulation
-Our problem setup is illustrated in Fig.
-2.
-Given a video of $N$ frames, let $M\in\mathbb{R}^{N\times P\times 3}$ be a pre-extracted point tracks tensor (Fig.
-2, left side).
-This tensor represents the two-dimensional information about a set of $P$ world points that are tracked throughout the video.
-Each element in the tensor, $M_{i,j}$, stores three values: $(x,y,o)$ where $x,y\in\mathbb{R}$ are respectively the observed horizontal and vertical locations of point $j$ in frame $i$, and $o\in\{0,1\}$ indicates whether point $j$ is observed in frame $i$ or not.
-Our goal is to learn a function $f$ to map the input point tracks tensor $M$ into a set of per-frame camera poses $\{R_{i},\mathbf{t}_{i}\}_{i=1}^{N}$ and per-frame 3D points $\{X_{i}\}_{i=1}^{N}$, where $R_{i}\in SO(3),\mathbf{t}_{i}\in\mathbb{R}^{3},X_{i}\in\mathbb{R}^{P\times 3}$ (Fig.
-2, right side).
-Discussion.
-We note that given enough point tracks, $M$ contains the 2D motion constraints that can be provided by the video.
-The source of each 2D motion constraint can be either from a global camera 3D motion or from an independent object motion.
-While the pixel appearance and semantic information could be used as well as inputs in our task, we find that the 2D motion signal $M$ gives useful geometric constraints for our setup.
-Moreover, this abstraction of the video helps our network generalize well to out-of-distribution videos with object classes that have not been observed during training but share similar 2D motion characteristics as we show in our experiments.
-Therefore $f$ only uses $M$ as input and does not rely on the input video once $M$ has been extracted.
-Figure 2: Our pipeline Our network takes as input a set of 2D point tracks (left) and uses several multi-head attention layers while alternating between the time dimension and the track dimension (middle).
+At inference time, our network predicts the dynamic structure and camera motion in a single feed-forward pass.
+Our network takes as input a set of 2D point tracks (left) and uses several multi-head attention layers while alternating between the time dimension and the track dimension (middle).
 The network predicts cameras, per-frame 3D points, and per-world point movement value (right).
-The 3D point colors illustrate the movement level values, such that points with high/low motion values are presented in blue/purple colors respectively.
+The 3D point internal colors illustrate the predicted 3D movement level values, such that points with high/low 3D motion are presented in red/purple colors respectively.
 These outputs are used to reproject the predicted points into the frames for calculating the reprojection error losses.
 See details in the text.
+The reader is encouraged to watch the supplementary video visualizations.
+Our approach.
+We propose TracksTo4D,1114D since we have three Euclidean coordinates with an additional time coordinate a novel approach for fast reconstruction of sparse dynamic 3D point clouds and camera poses from casual videos.
+Our main idea is to train a neural network on multiple videos to learn the mapping from the input image sequence to a sequence of the scene’s 3D point clouds and camera poses.
+After training, the trained network can be efficiently applied to new image sequences using a single feed-forward pass, avoiding costly optimization.
+To enhance the method’s ability to generalize across different types of videos and scenes, we made a crucial design choice: our approach processes point track tensors as input, rather than operating directly on the image sequence.
+Specifically, Each entry $(n,p)$ in these tensors represents the 2D position of a tracked point $p$ in a specific video frame $n$ [6].
+Our main insight is that point track tensors may exhibit more common motion patterns across casual video domains compared to image pixels.
+In other words, we argue that processing the raw point track data rather than scene-specific pixels or features may enable learning class and scene-agnostic internal feature representations for improved generalization.
+Importantly, recent advances in point tracking [10, 18] enable efficiently inferring these point tracks from casual videos using pre-trained models.
+These two properties make point track matrices an attractive input for our learning method.
+Following this design choice, we design our architecture according to two principles: (1) process point track tensors, which have a unique structure, and (2) encode meaningful prior knowledge about the reconstruction problem, as the problem is ill-posed in general.
+In the following, we address these desired properties.
+First, we design a network architecture that can effectively and efficiently handle point track inputs.
+To do that, we propose a novel layer design that takes into account the symmetries of the problem: the mapping we aim to learn, from point track matrices to 3D point clouds and camera poses, preserves two natural symmetries: (i) the points being tracked can be arbitrarily permuted without affecting the problem; (ii) the frames containing these points exhibit temporal structure, adhering to an approximate time-translation symmetry.
+Following the Geometric Deep Learning paradigm [7], we build upon recent theoretical advances in equivariant learning [27] and integrate these two symmetries into our network architecture using dedicated attention and positional encoding mechanisms.
+Second, a key challenge in predicting 3D dynamic motion and camera poses from 2D point tracks is that this problem is inherently ill-posed without additional constraints [3].
+To address this, we integrate a low-rank movement assumption into our architecture, following the seminal work of [6] which constrained output point clouds to be linear combinations of basis elements.
+Specifically, given an input point track tensor, our architecture equivariantly predicts a small set of input-specific basis elements.
+The output point clouds at each time frame are then defined as a linear combination of these basis elements, with the coefficients also predicted by the network.
+Notably, the first basis is assumed to fully represent the 3D static points in the video, while the remaining basis elements capture the 3D dynamic deviations.
+This structure effectively restricts the predicted point clouds to have a more specific form, making the problem more constrained.
+Our approach is trained on a dataset of extracted point track matrices [18] from raw videos without any 3D supervision by simply minimizing the reprojection errors, aiming to predict output point clouds that, after undergoing a perspective projection, will return the original 2D point tracks.
+In our experiments, TracksTo4D is trained on the Common Pets dataset [40].
+We evaluate our method on test data with GT camera poses and GT depth information for point tracks, and demonstrate that it produces comparable results to state-of-the-art methods, while having a significantly shorter inference time by up to $95\%$.
+In addition, we show the method’s ability to generalize to out-of-domain videos.
+Contributions.
+In summary, our contributions are (1) A novel modeling of the dynamic reconstruction problem via learning on point tracks without 3D supervision; (2) A novel deep learning architecture incorporating two key principles: accounting for the symmetry of the data and encoding low-rank structure in the predicted point clouds (3) Experiments demonstrating extremely fast inference time compared to baselines, accurate results, and strong generalization across other categories.
 2 Method
-4.1 Overview
+Problem formulation.
+Given a video of $N$ frames, let $M\in\mathbb{R}^{N\times P\times 3}$ be a pre-extracted $2D$ point tracks tensor (Fig.
+1, left side).
+This tensor represents the two-dimensional information about a set of $P$ world points that are tracked throughout the video.
+Each element in the tensor, $M_{i,j,:}$, stores three values: $(x,y,o)$ where $x,y\in\mathbb{R}$ are respectively the observed horizontal and vertical locations of point $j$ in frame $i$, and $o\in\{0,1\}$ indicates whether point $j$ is observed in frame $i$ or not.
+Our goal is to train a deep neural network to map the input point tracks tensor $M$ into a set of per-frame camera poses $\{R_{i}(M),\mathbf{t}_{i}(M)\}_{i=1}^{N}$ and per-frame 3D points $\{X_{i}(M)\}_{i=1}^{N}$, where $R_{i}(M)\in\mathbb{SO}(3),\mathbf{t}_{i}(M)\in\mathbb{R}^{3},X_{i}(M)\in\mathbb{R}^{P\times 3}$ (Fig.
+1, right side).
+Overview of our approach.
+Our method receives $M\in\mathbb{R}^{N\times P\times 3}$ as input.
+This tensor is being processed by a neural architecture composed of multi-head attention layers where the attention is applied in an alternating fashion on the $P$ and the $N$ dimensions in each layer.
+These layers are defined in Sec.
+2.1.
+After a composition of several such layers, the network uses the resulting features in $\mathbb{R}^{N\times P\times d}$ to predict $N$ camera poses in $\mathbb{SO}^{3}\times\mathbb{R}^{3}$ and $N$ point clouds in $\mathbb{R}^{N\times P\times 3}$.
+These $N$ point clouds are parameterized as a linear combination of $K$ input specific point cloud bases $B_{1}(M),\dots B_{K}(M)\in\mathbb{R}^{P\times 3}$.
+This is discussed in detail in Sec.
+2.2.
+Our network is trained in an unsupervised way on a dataset of videos by minimizing the reprojection error and other regularization losses (Sec.
+2.3) that are used to update the model parameters.
 Our pipeline is illustrated in Fig.
-2.
-Our method receives $M\in\mathbb{R}^{N\times P\times 3}$ as input.
-This tensor is being processed by a neural architecture composed of multi-head attention layers where the attention is calculated sequentially on the $P$ and the $N$ dimensions in each layer (Sec.
-4.3).
-After several such layers, we perform a projection to $N$ per-frame parameters and $P$ per-world-point parameters which determine the camera poses and 3D dynamic structure (Sec.
-4.2).
-This projection is done by first pooling over the secondary dimension and applying suitable layers to each one of the outputs to get the per-camera and per-world points parameters respectively.
-Our network is trained by minimizing the reprojection error and other regularization losses (Sec.
-4.4) that are used to update the model parameters.
-4.2 Constraining the Non-Rigid Structure and Motion
-Given the input 2D tracks, we wish to describe their motion by global camera motion and non-rigid self-motion.
-The 2D motion that is produced from rigid points constrains well the camera motion, but predicting the camera poses and the 3D non-rigid motion from 2D motion alone is an ill-posed problem without any constraints [3].
-We tackle this challenge in two ways.
-First, as done in classical Non-Rigid Structure from Motion [6], we constrain the output points to be formulated by a linear combination of basis elements.
-This constrains the 3D motion to be simple and prevents "jumps" in the unconstrained depth.
-More formally we assume that the non-rigid structure is formulated by a few basis elements $B_{1},\dots,B_{K}\in\mathbb{R}^{P\times 3}$ such that $X_{i}=\sum_{k=1}^{K}c_{ik}B_{k}$ where $\{c_{ik}\}_{k=1}^{K}$ are the $K$ coefficients which define the 3D structure of frame $i$.
-Secondly, we define the first basis element $B_{1}\in\mathbb{R}^{P\times 3}$ to be a rigid approximation of the points, such that the first basis coefficient in every frame is $1$, i.e.
-$c_{i1}=1$ for $i=1,\dots,N$.
-This helps in constraining the camera motion from the 2D motion of the rigid scene part.
-However, since this global rigid approximation cannot explain the 2D motion that is produced by the non-rigid world points, we model the reprojection error of the rigid approximation with $P$ zero-mean Cauchy distributions.
-For a non-rigid world point, $B_{1}$ by itself cannot produce low reprojection errors, and therefore its associated $\gamma$ is expected to be higher.
-The parameters of the Cauchy distributions, $\gamma_{1},\dots\gamma_{P}\in\mathbb{R}_{+}$ (Fig.
-2, "Movement Level Values") are predicted as well by the network and our training maximizes the likelihood of this assumption (see details in Sec.
-4.4).
-This allows our network to benefit from the rigid part for camera pose prediction while still being robust to the non-rigid motion.
-The rest of the basis elements determine the non-rigid deviation of the non-rigid points from the rigid approximation.
-In summary, our function $f$ gets $M\in\mathbb{R}^{N\times P\times 3}$ as an input and predicts per-frame outputs:
-$R_{i}\in SO(3),\mathbf{t}_{i}\in\mathbb{R}^{3},\{c_{ik}\}_{k=2}^{K},\text{for}\ \ i=1,\dots,N$ (1)
-and per-point outputs:
-$B_{1},\dots,B_{K}\in\mathbb{R}^{P\times 3},\gamma_{1},\dots,\gamma_{P}\in\mathbb{R}_{+}$ (2)
-which together define the non-rigid structure and motion.
-4.3 Architecture
+1
+2.1 Equivariant layers for point track tensors
 Following the geometric deep learning paradigm, our goal is to design a neural architecture that respects the underlying symmetries and structure of the data.
@@
 Our input is a tensor $M\in\mathbb{R}^{N\times P\times 3}$ representing a sequence of $N$ frames each with $P$ point coordinates.
-This structure gives rise to two key symmetries: First, the order of the $P$ points within each frame does not matter - permuting this axis results in an equivalent problem [27].
+This structure gives rise to two key symmetries: First, the order of the $P$ points within each frame does not matter - in other words, permuting this axis results in an equivalent problem [27].
 Formally, this axis has a permutation symmetry $S_{P}$ where $S_{P}$ is the symmetric group on $P$ elements.
 Second, along the temporal $N$ axis, we have an approximate translation symmetry arising from the ordered video sequence.
-This means that translating the time frames is required to result in nearly the same overall representation.
+This means that shifting the time frames is required to result in the same shift in our output.
 We model this with a cyclic group $C_{N}$ of order $N$.
 Both symmetries are illustrated in Fig.
-3.
+2.
+We note that while the cyclic group assumption may not be entirely accurate, we still find it useful as it helps us to derive appropriate parametric layers for our data, similar to how the convolutional layer is derived for data with translational symmetries such as images.
 Figure 2: The symmetry structure of our problem. Frames (vertical) have time translation symmetry while points (horizontal) have set permutation symmetry.
-Taken together, the full symmetry group of the input space is the direct product $\mathcal{G}=C_{N}\times S_{P}$ combining these frame and point permutation symmetries  [^1] This is different from the symmetry group studied in [29], where the temporal structure was not exploited.
-Next, we will design an architecture equivariant to $\mathcal{G}$, so that we can incorporate these domain priors and ensure that the model takes into account the underlying data structure.
-Network Architecture.
-Building on the analysis from the Deep Sets for Symmetric elements (DSS) approach [27], if we assumed we can use only linear layers, our architecture would take the form:
+Taken together, the full symmetry group of the input space is the direct product $\mathcal{G}=C_{N}\times S_{P}$ combining these time and point permutation symmetries, acting on $\mathbb{R}^{N\times P\times 3}$ by $((t,\sigma)\cdot M)_{n,p,j}=M_{t^{-1}(n),\sigma^{-1}(p),j}$ for $(t,\sigma)\in\mathcal{G}$ [^2] This is different from the symmetry group studied in [29], where the temporal structure was not exploited..
+Next, we will design an architecture equivariant to $\mathcal{G}$, to ensure that the model takes into account the symmetries above.
+Linear equivariant layers.
+Point track tensors can be viewed as a collection of $N$ individual point tracks, each of which exhibits translational symmetry.
+The scenario where an object comprises a set of elements with their own symmetry group, such as a set of images or graphs, was previously explored in [27].
+In that work, the authors characterized the general linear equivariant layer structure in such cases, termed the Deep Sets for Symmetric Elements (DSS) layer.
+Building on the DSS approach, our basic linear equivariant layer for the point track tensors $M$ would take the form:
 $F(M)_{:,j}=L_{1}(M_{:,j})+\sum_{j^{\prime}=1}^{P}L_{2}(M_{:,j^{\prime}})$ (1)
-where $L$ is a translation equivariant function, $M_{:,j}\in\mathbb{R}^{N\times d}$ are the columns of $M$ representing all the inputs for a specific tracked point, $F(M)_{:,j}\in\mathbb{R}^{N\times d^{\prime}}$ is the output column and we have $d,d^{\prime}$ input and output feature channels respectively.
-While valid, this linear layer design might not be the optimal choice.
-To enhance the model, we incorporate additional nonlinear layers that are well-suited for sequence and set structures data.
-One particularly effective choice is the transformer layer [46].
-Importantly, by leveraging attention mechanisms and temporal positional encodings, the transformer processes sequential input while respecting its ordering and context.
-Specifically, our layer $F$ is formulated very similarly to Equation (3) and is implemented by replacing the convolutions ($L_{i}$) and summations in Equation (3) with two self-attention mechanisms and suitable temporal positional encoding across the $N$ dimension.
-We provide more architecture and implementation details in the supplementary.
-4.4 Training
-Given a dataset of point track representations: $\{M_{1},\dots,M_{L}\}$, our method trains the function $f$ to minimize the loss function over all elements in the dataset:
-[…]
-Our loss function is defined by:
-$\mathcal{L}=\lambda_{\text{Reproject}}\mathcal{L}_{\text{Reproject}}+\lambda_{\text{Rigid}}\mathcal{L}_{\text{Rigid}}+\lambda_{\text{Negative}}\mathcal{L}_{\text{Negative}}+\lambda_{\text{Sparse}}\mathcal{L}_{\text{Sparse}}$ (5)
-The first loss is the reprojection error given by the structure and the cameras predicted by $f$:
+where $L_{i}$ are linear translation equivariant function (i.e.
+convolutions), $M_{:,j}\in\mathbb{R}^{N\times d}$ are the columns of $M$ representing all the inputs for a specific tracked point, $F(M)_{:,j}\in\mathbb{R}^{N\times d^{\prime}}$ is the output column and $d,d^{\prime}$ are the input and output feature channels respectively.
+To construct a neural network, these layers can be interleaved with pointwise nonlinearities, similar to basic convolutional neural networks.
+Implementation via transformers and positional encoding.
+While the linear layer design is reasonable, it may not be the optimal choice.
+To enhance the model, we design a new layer whose structure follows Equation (1), but incorporates nonlinear layers in the form of transformers [47].
+Specifically, our layer $F$ is formulated similarly to Equation (1), but instead of convolutions ($L_{i}$) and summations, it utilizes two self-attention mechanisms and suitable temporal positional encoding across the $N$ dimension.
+Formally, our basic layer $F:\mathbb{R}^{N\times P\times d}\rightarrow\mathbb{R}^{N\times P\times d^{\prime}}$ is computed via four steps, which are described below:
+$\mathbf{\bar{q}}_{ij}=\bar{W}^{Q}M_{ij},~{}\mathbf{\bar{k}}_{ij}=\bar{W}^{K}M_{ij},~{}\mathbf{\bar{v}}_{ij}=\bar{W}^{V}M_{ij}\Rightarrow\bar{M}_{ij}=\sum_{i^{\prime}=1}^{N}\frac{\exp(\mathbf{\bar{q}}_{ij}\cdot\mathbf{\bar{k}}_{i^{\prime}j})}{\sum_{l=1}^{N}\exp(\mathbf{\bar{q}}_{ij}\cdot\mathbf{\bar{k}}_{lj})}\mathbf{\bar{v}}_{i^{\prime}j}$ (2)
+$\mathbf{q}_{ij}=W^{Q}\bar{M}_{ij},~{}\mathbf{k}_{ij}=W^{K}\bar{M}_{ij},~{}\mathbf{v}_{ij}=W^{V}\bar{M}_{ij}\Rightarrow F(M)_{ij}=\sum_{j^{\prime}=1}^{P}\frac{\exp(\mathbf{q}_{ij}\cdot\mathbf{k}_{ij^{\prime}})}{\sum_{l=1}^{P}\exp(\mathbf{q}_{ij}\cdot\mathbf{k}_{il})}\mathbf{v}_{ij^{\prime}}$ (3)
+Here, $M_{i,j}\in\mathbb{R}^{d}$ are the features associated with the $j$-th point in the $i$-th frame.
+The attention mechanism defined in the first equation above (2) is augmented with standard temporal positional encoding in the first layer and replaces the translation equivariant function $L_{i}$ applied to the columns of $M$ (Eq.(1)).
+The attention in the second equation (3) implements the set aggregation (summation) (also in Eq.(1)) applied to the rows of $M$.
+As commonly done, we use transformers with 16 attention heads [47].
+2.2 Constraining 3D motion and camera poses via low-rank assumption
+Given our 2D tracks, we aim to characterize the motion of the points by decomposing them into the global camera motion and the 3D motion of objects in the scene.
+The 2D motion of static scene points provides useful constraints for estimating the camera motion.
+However, as previously mentioned, predicting camera and dynamic 3D motion solely from 2D motion is an ill-posed problem without additional constraints [3].
+We tackle this challenge by adding two mechanisms to our architecture: (1) low-rank movement assumption; and (2) specific modeling of the static scene for camera estimation.
+Low-rank movement assumption.
+First, motivated by classical orthographic Non-Rigid Structure from Motion [6], we constrain the output points to be formulated by a linear combination of input-specific basis elements.
+Specifically, given the input 2D point tracks, $M\in\mathbb{R}^{N\times P\times 3}$, our network predicts $K$ point clouds: $B_{1}(M),\dots,B_{K}(M)\in\mathbb{R}^{P\times 3}$ and $N(K-1)$ linear coefficients, $\{c_{1k}(M)\}_{k=2}^{K},\dots\{c_{Nk(M)}\}_{k=2}^{K}$ such that $X_{i}(M)=B_{1}(M)+\sum_{k=2}^{K}c_{ik}(M)B_{k}(M),$ where $X_{i}(M)\in\mathbb{R}^{P\times 3}$ is the 3D point cloud at frame $i$.
+The point clouds and coefficients are computed by taking the output of the last equivariant layer as defined in the previous section and applying invariant aggregations on the respective dimension resulting in equivariant and invariant outputs.
+See more details in the appendix.
+We note that we deliberately chose the coefficient of $B_{1}(M)$ to be the constant $1$, the reason is explained in the next paragraph.
+Specific modeling of the static scene for camera estimation.
+Frequently, casual video data of dynamic scenes contains many static regions, which can be used to determine camera poses [61].
+We leverage this observation by treating the first basis element $B_{1}(M)\in\mathbb{R}^{P\times 3}$ as a static approximation for all scene points and encourage $B_{1}(M)$ as well as the output camera poses to explain the 2D observations according to this approximation using a "static" reprojection loss ($\mathcal{L}_{\text{Static}}$, defined in the next section).
+We note, however, that a static point cloud is not likely to produce low reprojection errors for the non-static components, thus the reprojection error necessitates robustness to substantial errors from the non-static elements.
+To address this, our network predicts (equivariantly) $P$ motion level values $\gamma_{1}(M),\dots,\gamma_{P}(M)\in\mathbb{R}_{+}$, one for each point in our dynamic point cloud, which we use to weight the reprojection errors from $B_{1}(M)$.
+The main idea is to give less weight to non-static points so that the static projection loss can disregard them.
+Specifically, inspired by [59], each $\gamma_{i}(M)$ defines a Cauchy distribution that models the reprojection errors for its associated world point, such that a world point with higher $\gamma$ is expected to produce a wider error distribution.
+Empirically, as noted by [59], the Cauchy distribution tends to be more robust for modeling reprojection error uncertainties compared to Gaussian noise modeling [19].
+Then, $\mathcal{L}_{\text{Static}}$, minimizes the negative log-likelihood under this assumption.
+See details in Sec.
+2.3.
+2.3 Training and losses
+Model outputs.
+Given the input 2D point tracks $M\in\mathbb{R}^{N\times P\times 3}$, our network produces outputs as a function of $M$: linear bases and coefficients $B_{1}(M),\dots,B_{K}(M)\in\mathbb{R}^{P\times 3},\{c_{1k}(M)\}_{k=2}^{K},\dots,\{c_{Nk}(M)\}_{k=2}^{K}\in\mathbb{R}$ which define a dynamic point cloud $X_{1}(M),\cdots,X_{N}(M)\in\mathbb{R}^{P\times 3}$,$\gamma_{1}(M),\dots,\gamma_{P}(M)\in\mathbb{R}_{+}$ movement level values, and $(R_{1}(M),\mathbf{t}_{1}(M)),\dots,(R_{N}(M),\mathbf{t}_{N}(M))\in SO(3)\times\mathbb{R}^{3}$ camera poses.
+We use these network outputs to define an self-supervised loss function with respect to the current network weights and $M$ which is defined by:
+$\mathcal{L}=\lambda_{\text{Reproject}}\mathcal{L}_{\text{Reproject}}+\lambda_{\text{Static}}\mathcal{L}_{\text{Static}}+\lambda_{\text{Negative}}\mathcal{L}_{\text{Negative}}+\lambda_{\text{Sparse}}\mathcal{L}_{\text{Sparse}}$ (4)
+Reprojection loss.
+The reprojection loss encourages the consistency between the output 3D point clouds and camera poses, to the input 2D observations:
 $\mathcal{L}_{\text{Reproject}}=\frac{1}{\sum_{i=1}^{N}\sum_{j=1}^{P}M_{ij}^{o}}\sum_{i=1}^{N}\sum_{j=1}^{P}M_{ij}^{o}\mathcal{R}(X_{ij},R_{i},\mathbf{t}_{i},M_{ij}^{xy})$ (5)
@@
 $\mathcal{R}(X_{ij},R_{i},\mathbf{t}_{i},M_{ij}^{xy})=\left\lVert\frac{(R_{i}^{T}(\mathbf{X}_{ij}-\mathbf{t}_{i}))_{1,2}}{(R_{i}^{T}(\mathbf{X}_{ij}-\mathbf{t}_{i}))_{3}}-M_{ij}^{xy}\right\rVert$ (6)
+Static loss.
 As discussed in Sec.
-4.2, the first predicted basis element $B_{1}\in\mathbb{R}^{P\times 3}$ defines a rigid (fixed in time) approximation.
-Our network also predicts a movement coefficient $\gamma_{j}$ for each world point that defines a zero-mean Cauchy distribution.
+2.2, to better constrain the camera poses, the first predicted basis element $B_{1}(M)\in\mathbb{R}^{P\times 3}$ defines a static (fixed in time) point cloud approximation.
+Our network also predicts a movement coefficient $\gamma_{j}(M)$ for each world point that defines a zero-mean Cauchy distribution.
 Given $\gamma_{j}$ and the reprojection error $r_{ij}=\mathcal{R}(B_{1j},R_{i},\mathbf{t}_{i},M_{ij}^{xy})$ of the $j^{th}$ point of $B_{1}$ that is projected by the $i^{th}$ camera, the negative log-likelihood of $r_{ij}$ distributed according to the $\gamma_{j}$-zero-mean Cauchy distribution is proportional to:
 $\mathcal{C}(r_{ij},\gamma_{j})=\log\left(\gamma_{j}+\frac{r_{ij}^{2}}{\gamma_{j}}\right)$ (7)
-Our rigid loss calculates the mean negative log-likelihood over all observed points in all frames:
-$\mathcal{L}_{\text{Rigid}}=\frac{1}{\sum_{i=1}^{N}\sum_{j=1}^{P}M_{ij}^{o}}\sum_{i=1}^{N}\sum_{j=1}^{P}M_{ij}^{o}\mathcal{C}(\mathcal{R}(B_{1j},R_{i},\mathbf{t}_{i},M_{ij}^{xy}),\gamma_{j})$ (9)
-As in [29] we regularize the points to be in front of the camera by:
+Note, that this loss reduces the contribution of the reprojection errors for points with high $\gamma$, but also encourages $\gamma$ to be small, i.e.
+encouraging the static point cloud to represent the dynamic scene when possible.
+Our static loss is the mean negative log-likelihood over all observed points in all frames:
+$\mathcal{L}_{\text{Static}}=\frac{1}{\sum_{i=1}^{N}\sum_{j=1}^{P}M_{ij}^{o}}\sum_{i=1}^{N}\sum_{j=1}^{P}M_{ij}^{o}\mathcal{C}(\mathcal{R}(B_{1j},R_{i},\mathbf{t}_{i},M_{ij}^{xy}),\gamma_{j})$ (8)
+Regularization losses.
+As in [29] we regularize the observed points to be in front of the camera by:
 $\mathcal{L}_{\text{Negative}}=-\sum_{i=1}^{N}\sum_{j=1}^{P}M_{ij}^{o}\text{ Min}(R_{i}^{T}(\mathbf{X}_{ij}-\mathbf{t}_{i}))_{3},0)$ (9)
-We further find it beneficial to regularize the deviation from the rigid approximation $B_{1}$ to be sparse at locations with low $\gamma$ values:
+We further find it beneficial to regularize the deviation from the static approximation $B_{1}$ to be sparse for static points, i.e. points with low $\gamma$ values:
 $\mathcal{L}_{\text{Sparse}}=\frac{1}{P(K-1)}\sum_{k=2}^{K}\sum_{j=1}^{P}\frac{1}{3\gamma_{j}}\left(|B_{kj1}|+|B_{kj2}|+|B_{kj3}|\right)$ (10)
 where $\gamma$ is detached from the gradient calculation for this loss.
-ATE $\downarrow$ RPE Trans $\downarrow$ RPE Rot $\downarrow$ Inference $\downarrow$
-(mm) (mm) (deg) Time
-DROID-SLAM [43] 5.08 3.60 0.20 10 seconds
-ParticleSfM [64] 12.79 6.95 0.51 11 minutes
-RCVD[21] 43.95 25.77 2.31 20 minutes
-CasualSAM [62] 6.90 3.95 0.22 130 minutes
-Ours (cats only) 8.96 3.79 0.23 0.16(+8.6) seconds
-Ours (cats only)+BA 4.22 2.86 0.17 0.4(+8.6) seconds
-Ours (dogs only) 8.03 3.54 0.23 0.16(+8.6) seconds
-Ours (dogs only)+BA 4.19 2.83 0.17 0.4(+8.6) seconds
-Ours (dogs+cats) 8.11 3.68 0.24 0.16(+8.6) seconds
-Ours (dogs+cats)+BA 4.21 2.86 0.17 0.4(+8.6) seconds
-Table 1: Camera Poses Accuracy.
-Quantitative comparisons between our method to previous methods for camera accuracy and running time.
-For our method, we compare three versions of (1) training only on cats, (2) training only on dogs, and (3) training on both.
-We evaluate all methods on an unseen test set of cats and dogs.
-We further show our accuracies after one round of Bundle Adjustment.
-For our method’s running time, we also mention the point tracks extraction time in parenthesis (+8.6 seconds) which is performed by [18] as a preprocess.
-Abs Rel $\downarrow$ $\delta<1.25\uparrow$ $\delta<1.25^{2}\uparrow$ $\delta<1.25^{3}\uparrow$
-Dynamic All Dynamic All Dynamic All Dynamic All
-RCVD[21] 0.40 3.6e+7 0.43 0.72 0.75 0.90 0.92 0.96
-MiDaS [4] 0.16 6.1e+4 0.78 0.71 0.97 0.88 1.00 0.93
-CasualSAM [62] 0.09 0.06 0.93 0.97 0.99 0.99 1.00 1.00
-Our (cats only) 0.11 0.08 0.88 0.92 0.99 0.98 1.00 1.00
-Our (dogs only) 0.12 0.08 0.85 0.91 0.99 0.99 1.00 1.00
-Our (dogs+cats) 0.12 0.08 0.85 0.91 0.98 0.98 1.00 1.00
-Table 2: Depth Accuracy We show a comparison to previous methods on the predicted depth for the trajectories compared to their GT depths.
-We compare three instances of our method, each trained on a different partition of the dataset.
+Table 1: Pet evaluation.
+Top: Baseline method results for structure or camera estimation (or both).
+Bottom: Our results with several configurations.
+(C),(D), or (CD) respectively indicate the object categories in the training set: cats, dogs, or both.
+BA and FT respectively indicate a post-processing of Bundle Adjustment or fine-tuning.
+Abs Rel $\downarrow$ $\delta<1.25\uparrow$ $\delta<1.25^{2}\uparrow$ $\delta<1.25^{3}\uparrow$ ATE $\downarrow$ RPE Trans $\downarrow$ RPE Rot $\downarrow$ Time
+Dyn. All Dyn. All Dyn. All Dyn. All (mm) (mm) (deg) (min)
+D-SLAM [44] - - - - - - - - 5.08 3.60 0.20 0.16
+ParticleSFM [61] - - - - - - - - 12.79 6.95 0.51 11.00
+RCVD [22] 0.40 3.6E+07 0.43 0.72 0.75 0.90 0.92 0.96 43.95 25.77 2.31 20.00
+CasualSAM [59] 0.09 0.06 0.93 0.97 0.99 0.99 1.00 1.00 6.90 3.95 0.22 1.3E+02
+MiDaS [4] 0.16 6.2E+04 0.78 0.71 0.97 0.88 1.00 0.93 - - - 0.15
+Ours (C) 0.11 0.08 0.88 0.92 0.99 0.98 1.00 1.00 8.96 3.79 0.23 0.15
+Ours (C)+BA 0.11 0.08 0.88 0.92 0.99 0.98 1.00 1.00 4.22 2.86 0.17 0.15
+Ours (C)+FT 0.09 0.06 0.90 0.96 1.00 0.99 1.00 1.00 4.00 2.74 0.16 4.86
+Ours (D) 0.12 0.08 0.85 0.91 0.99 0.99 1.00 1.00 8.03 3.54 0.23 0.15
+Ours (D)+BA 0.12 0.08 0.85 0.91 0.99 0.99 1.00 1.00 4.19 2.83 0.17 0.15
+Ours (D)+FT 0.09 0.06 0.88 0.96 1.00 0.99 1.00 1.00 3.98 2.74 0.16 4.86
+Ours (CD) 0.12 0.08 0.85 0.91 0.98 0.98 1.00 1.00 8.11 3.68 0.24 0.15
+Ours (CD)+BA 0.12 0.08 0.85 0.91 0.98 0.98 1.00 1.00 4.21 2.86 0.17 0.15
+Ours (CD)+FT 0.09 0.06 0.90 0.96 1.00 0.99 1.00 1.00 3.98 2.74 0.16 4.86
+Table 2: Out-of-training-domain evaluation . Evaluation metrics on monocular videos from [58]. The table has the same structure as Tab. 1.
+Abs Rel $\downarrow$ $\delta<1.25\uparrow$ $\delta<1.25^{2}\uparrow$ $\delta<1.25^{3}\uparrow$ ATE $\downarrow$ RPE Trans $\downarrow$ RPE Rot $\downarrow$ Time
+Dyn. All Dyn. All Dyn. All Dyn. All (mm) (mm) (deg) (min)
+D-SLAM [44] - - - - - - - - 7.96 10.91 0.07 0.18
+ParticleSFM [61] - - - - - - - - 26.66 23.83 0.20 2.13
+RCVD [22] 0.19 2.6E+05 0.69 0.75 0.95 0.95 0.96 0.98 1.6E+02 3.2E+02 3.43 7.00
+CasualSAM [59] 0.05 0.03 0.95 0.99 0.99 1.00 1.00 1.00 7.81 10.09 0.06 22.00
+MiDaS [4] 2.8E+04 2.7E+05 0.59 0.58 0.73 0.72 0.83 0.80 - - - 0.02
+Ours (C) 0.08 0.06 0.89 0.95 0.99 0.99 0.99 1.00 32.06 47.99 0.45 0.04
+Ours (C)+BA 0.08 0.06 0.89 0.95 0.99 0.99 0.99 1.00 8.67 12.36 0.08 0.04
+Ours (C)+FT 0.07 0.03 0.94 0.98 0.99 1.00 1.00 1.00 7.98 11.64 0.08 0.59
+Ours (D) 0.08 0.07 0.92 0.93 0.99 0.98 0.99 1.00 33.77 51.64 0.61 0.04
+Ours (D)+BA 0.08 0.07 0.92 0.93 0.99 0.98 0.99 1.00 8.40 12.06 0.08 0.04
+Ours (D)+FT 0.05 0.03 0.97 0.99 0.99 1.00 0.99 1.00 8.15 11.88 0.09 0.59
+Ours (CD) 0.10 0.08 0.93 0.94 0.99 0.99 1.00 1.00 36.17 53.94 0.67 0.04
+Ours (CD)+BA 0.10 0.08 0.93 0.94 0.99 0.99 1.00 1.00 8.62 12.49 0.08 0.04
+Ours (CD)+FT 0.06 0.03 0.97 0.99 0.99 1.00 0.99 1.00 8.04 11.84 0.09 0.59
+Table 3: Ablation study. The contribution of different parts from our method. See details in the text.
 Abs Rel $\downarrow$ $\delta<1.25\uparrow$ $\delta<1.25^{2}\uparrow$ $\delta<1.25^{3}\uparrow$ Rep.(pix.) $\downarrow$ ATE $\downarrow$ RPE Trans $\downarrow$ RPE Rot $\downarrow$
@@
 Set of Sets 0.27 0.15 0.60 0.77 0.87 0.94 0.97 0.99 9.86 5.33 16.87 5.53 0.39
-No $\mathcal{L}_{\text{Rigid}}$ 0.77 0.36 0.25 0.46 0.48 0.70 0.68 0.82 1.00 0.86 96.20 29.86 0.99
+No $\mathcal{L}_{\text{Static}}$ 0.77 0.36 0.25 0.46 0.48 0.70 0.68 0.82 1.00 0.86 96.20 29.86 0.99
 No $\gamma$ 0.22 0.16 0.66 0.73 0.93 0.91 0.99 0.97 4.54 2.41 13.91 4.86 0.29
@@
 Full 0.11 0.08 0.88 0.92 0.99 0.98 1.00 1.00 3.98 1.97 8.96 3.79 0.23
-Table 3: Ablation Study. We evaluate the contribution of different parts from our method by changing one property each time. See details in the text.
-Figure 4: Qualitative Results.
-We present qualitative results in the same format as in Fig.
-1.
-Top.
-3 frames from 2 different test video sequences with point tracks marked with corresponding colors.
-Bottom.
-A 3D visualization of the outputs of our method, from two time stamps.
-As the scene is observed from the same viewpoint in each pair, a non-rigid movement can be observed.
 3 Experiments
-5.1 Training Details
+In this section, we conduct experiments to verify our proposed network’s performance on real-world casual videos.
+We began by training the network on specific domains and then evaluated its accuracy and running time on unseen videos from both, training and unseen domains.
+Training procedure.
 We trained our network on the cat and dog partitions from the COP3D dataset [40], which contains a diverse set of casual real-world videos of pets.
-We trained our network from scratch three times to test our generalization capability between semantic categories: once on the cat partition, once on the dog partition, and once on both partitions combined.
-In total, we used $733$ cat videos and $753$ dog videos for training.
-We trained our networks for $7000$ and $3500$ epochs for the single-class and multi-class setups respectively.
-Training our method lasts about one week on a single Tesla V100 GPU with 32GB memory.
-We used the Adam optimizer [20] with a learning rate of $10^{-4}$.
-For our non-rigid bases $B_{1},...,B_{K}$ (Sec.
-4.2) we used $K=12$.
-Our method assumes known camera internal parameters which are provided by the dataset and used to normalize the point tracks as a preprocess.
-5.2 Evaluation Data
-While the COP3D dataset provides cameras that were extracted by COLMAP [37], we note that the non-rigid structure was captured as well in part of their reconstruction which indicates that its reconstruction might not be accurate.
+Our networks were trained from scratch three times to test our generalization capability between semantic categories: once on the cat partition, once on the dog partition, and once on both partitions combined.
+Training technical details are provided in the appendix.
+We use $K=12$ bases in all our experiments (Sec.
+2.2).
+Evaluation data.
+To assess our framework’s performance on pet videos, we curated a new dataset333 While the COP3D dataset provides cameras that were extracted by COLMAP [38], we note that this evaluation data is insufficient in our case.
+This is because the dynamic structure was captured as well in part of their reconstruction which indicates that its reconstruction might not be accurate.
 Furthermore, the coordinates system units of these reconstructions are unknown.
-Finally, this dataset does not have any depth map information for evaluating the non-rigid structure.
-For evaluation, we used an RGBD sensor and collected 21 casual videos of dogs and cats, each with 50 frames.
+Finally, this dataset does not have any depth map information for evaluating the dynamic structure.
+consisting of 21 casual videos of dogs and cats, each video containing 50 frames.
+These videos were captured using an RGBD (RGB-Depth) sensor.
 The depth maps were used as ground truth for evaluating the reconstructed structure.
@@
 The cameras were scaled to millimeter units using the provided GT depth.
-We verified that this data includes enough non-rigid motion, by also including several videos that COLMAP failed to reconstruct without the masks.
-We further verified manually that the camera trajectories look reasonable.
-Note that this is test data that our network did not see during training and was not used to tune our hyperparameters.
-5.3 Results
-We evaluated our method on the test data.
+Note that our network did not see this test data during training and it was not used to tune our hyperparameters.
+Additionally, to evaluate our method on out-of-domain evaluation data, we used the Nvidia Dynamic Scenes Dataset [58].
+Specifically, while our network was trained on pet videos, this dataset contains other dynamic object types, e.g.
+human, balloon, truck, and umbrella, with a different camera motion type.
+The dataset contains 8 dynamic scenes which are captured by 12 synchronized cameras, enabling accurate depth estimation which is treated as GT for evaluating monocular depth estimation.
+The ground truth camera poses were calculated by [38] with the synchronized multiview camera rig and the ground truth dynamic masks.
+Similarly to [24] we simulated 8 monocular dynamic video sequences using the camera rig, each with 24 frames, and used them for evaluation.
+Evaluation results.
 Qualitative visualizations are presented in Fig.
-1 and Fig.
+3.444The reader is encouraged to watch the supplementary videos for better 4D perception.
+We also show a visualization of the movement level values, $\gamma$ in Fig.
 4.
-We also show a visualization of the movement level values, $\gamma$ in Fig.
-5.
-We compared our method to the state-of-the-art methods and evaluated both, the camera poses and the structure.
-The camera poses were evaluated using the Absolute Translation Error(ATE), the Relative Translation Error(RTE), and the Relative Rotation Error(RRE) metrics after coordinates system alignment.
-The results are presented in Tab.
-1.
-We compare three training possibilities of our method of training only on cats, only on dogs, and on both.
-As can be seen in the table, the results are similar in all 3 cases, even though the dest set contains samples of both dogs and cats.
-This shows that our network generalizes well to unseen animal classes.
+For comparisons, we chose state-of-the-art methods that as our method, can be applied to raw casual videos that were captured by standard pinhole camera models and do not need any static or dynamic segmentation.
+We evaluated both, the camera poses and the structure accuracies.
+Comparison results for the pet-test-set and out-of-domain dataset are presented in Tables 1 and 2 respectively.
+The camera poses are evaluated compared to the GT, using the Absolute Translation Error(ATE), the Relative Translation Error(RTE), and the Relative Rotation Error(RRE) metrics after coordinates system alignment.
+We compare three training configurations of our method of training only on cats, only on dogs, and on both.
+As can be seen in the tables, the results are similar in all 3 cases.
 Our output camera poses as inferred by the network are already accurate and outperform some of the prior methods.
-We further show the results of our method after a single and short round of Bundle Adjustment, which makes our method the most accurate one.
-Tab.
-1 also compares the running times between the methods.
-It can be seen that our method, even with bundle adjustment is the fastest one when most of the time (8.6 seconds) is used to extract the input point tracks with [18].
-To evaluate the structure we compare our method using depth evaluation metrics [13].
-The results are presented in Tab.
-2, and show that our inferred structure is comparable to the state-of-the-art [62] while taking significantly shorter running times (a few seconds for our method versus more than two hours for [62]).
-We also note that in contrast to other methods in the table, we are not using any supervision from a depth-from-single-image network.
+We further show the results of our method after a single and short round of Bundle Adjustment, which makes our method better than all baselines on the pet sequences, and comparable on the out-of-domain cases.
+Importantly, Tables 1 and 2 also compare the method’s running times.
+It can be seen that our method, even with bundle adjustment, is the fastest camera prediction method.
+Tables 1 and 2 also show structure evaluation with the depth evaluation metrics [11] on the sampled point tracks.
+They demonstrate that our inferred structure is almost comparable to the state-of-the-art [59] while taking significantly shorter running times (a few seconds for our method versus more than two hours for [59] on pet videos).
+Short (0.6-5 minutes), per-sequence fine-tuning makes our method’s accuracy comparable to [59], see appendix for details.
+In terms of running time, our method is a bit slower than MiDaS [4], which only provides depth maps without cameras, but achieves much better results.
+We note that in contrast to the other methods that predict the dynamic depth, ours does not use any depth-from-single-image prior.
+Figure 3: Qualitative Results.
+Top.
+Frames from 2 different test video sequences with point tracks marked with corresponding colors.
+Bottom.
+A 3D visualization of our method’s outputs, from two time stamps.
+The camera trajectory is present as gray frustums, whereas the current camera is marked in red.
+The reconstructed 3D scene points are presented in corresponding colors to the input tracks on the top.
+The scene is observed from the same viewpoint, enabling the visualization of the dynamic reconstructed structure.
 Figure 4: $\gamma$ Visualization.
 We show a visualization of the $\gamma$ outputs of our network that are described in Sec.
-4.2.
+2.2.
 In each video sequence, we show the input tracks, where each color visualizes its movement level value, $\gamma$.
-Purple marks rigid points with low $\gamma$ whereas red marks non-rigid points with high $\gamma$.
-Note, that our network did not get any direct supervision for these values, but the raw point tracks predictions from [18].
-To emphasize our generalization, the $\gamma$ visualizations for cats were produced by the model that was only trained on dogs and vice versa.
-5.4 Ablation Study
+Purple marks static points with low $\gamma$ whereas red marks dynamic points with high $\gamma$.
+Note, that our network did not get any direct supervision for these values, but only the raw point tracks predictions from [18].
+The $\gamma$ visualizations for cats were produced by the model that was only trained on dogs and vice versa.
+We note that our model generalizes well to out-of-domain (non-pet) cases as well.
+Ablation study
 To evaluate the contribution of our different method parts we run an ablation study which is presented in Tab.
 3.
-In this study, the training was always done on the cats partition from COP3D and evaluated on our test data which contains dogs and cats.
-More specifically, we ablate on our transformer architecture by taking the architecture suggested by [29] ("Set of Sets") or the DSS architecture [27] ("DSS").
+In this study, the training was always done on the cat partition from COP3D and evaluated on our test data which contains dogs and cats.
+First, we performed an ablation study on our transformer architecture by taking the architecture suggested by [29] ("Set of Sets") or the DSS architecture that uses only linear layers [27] ("DSS").
 As the table shows our architecture ("Full") achieved significantly better results.
-We further ignored the $\gamma$ outputs and used regular reprojection error on $B_{1}$ for all points ("No $\gamma$").
-We further ablate our sparsity loss ("No $\mathcal{L}_{\text{Sparse}}$") and the rigid loss ("No $\mathcal{L}_{\text{Rigid}}$").
-In both cases, the error increased whereas in the later one, the results became much worse.
-We further ablate the choice of $K=12$ as the number of non-rigid bases, by trying two extreme numbers of $K=30,K=2$ (we saw no significant differences when we used nearby choices such as $K=11$).
+To test the losses in our framework, we also evaluated the following: (1) ignoring the $\gamma$ outputs and using regular reprojection error on $B_{1}$ for all points ("No $\gamma$"); (2) removing our sparsity loss ("No $\mathcal{L}_{\text{Sparse}}$"); and (3) removing the static loss ("No $\mathcal{L}_{\text{Static}}$").
+In all cases, the error increased whereas in the later one, the results became much worse.
+We further ablate the choice of $K=12$ as the number of linear bases, by trying 2 extreme numbers of $K=30,K=2$ (we saw no significant differences when we used nearby choices such as $K=11$).
 As can be seen in the table, when $K=30$ the output is not regularized enough and produces a higher depth error for the dynamic part.
@@
 (pix.)") gets higher due to over-regularization.
-Overall the ablation study justifies our design choices ("Full").
-5.5 Limitations and Future Work
-Our method cannot handle videos with too rapid motion, and in general, is limited by the accuracy of the tracking method [18].
+Overall, this study justifies our design choices ("Full").
+4 Related Work
+Simultaneous Localization and Mapping (SLAM) and Structure from Motion (SfM)
+SfM pipelines seek to recover static 3D structure and camera poses from unordered images.[45, 41, 38, 1, 51].
+Learning-free pipelines [38] are effective but require repeated applications of Bundle Adjustment (BA) [46].
+[29, 8] presented a method for learning prior from a dataset of multiview image sets, to accelerate SfM pipelines by using equivariant deep networks.
+Monocular Simultaneous Localization and Mapping (SLAM) methods [30, 31, 12, 57, 5, 50, 62, 63, 42] extract camera poses from video sequences while defining a scene map with keyframes.
+These methods assume static scenes, fail to produce the cameras in scenes with large portions of dynamic motion, and cannot reproduce dynamic parts of the scene.
+DROID-SLAM [44] used synthetic data with ground truth 3D supervision for learning to predict camera poses via deep-based BA on keyframes while excluding dynamic objects.
+ParticleSfM [61] filters out 2D dynamic content for reproducing the cameras in dynamic scenes, using its pre-trained motion prediction network.
+Both, [44, 61] do not infer the dynamic 3D structure.
+Orthographic Non-Rigid SfM (NRSfM)
+Bregler et al.
+[6] introduced a factorization method for computing a non-rigid structure and rotation matrices from a point track matrix, by assuming a low dimensional basis model.
+While follow-up papers improved accuracies with different regularizations [23, 9, 33, 16] or neural representations [32, 21, 39], the orthographic camera model assumption is in general not valid for casual videos.
+Furthermore, these methods often assume background subtraction as a preprocessing.
+Even though a follow-up work proposed factorization solutions for pinhole cameras [14], its sensitivity to noise [17], makes it impractical for casual videos.
+Test-time optimization for dynamic scenes
+Recent methods [26, 22, 59, 60] finetuned the monocular depth estimation from a pre-trained model [37, 36] using optical flow constraints [43], for obtaining consistent dense depth maps for a monocular video.
+[59] further optimized motion maps for handling scenes with large dynamic motion.
+[52, 13] use depth from single image estimations, to improve novel view synthesis in dynamic scenes.
+[25] further optimizes for the unknown camera poses together with the dynamic radiance field optimization.
+[34, 35] model a single deformable surface from a monocular video by applying isometric constraints.
+LASR [54], ViSER [55] and BANMo [56] optimize for a dynamic surface by assuming rigid bones and linear blend skinning weights.
+However, all the above-mentioned methods require per-scene optimization, resulting in slow inference.
+Recently, [40] presented the Common Pets in 3D (COP3D) dataset that contains casual, in-the-wild videos of pets, and used it to learn priors for novel view synthesis in dynamic scenes.
+Point tracking
+There has been a recent advance in 2D point tracking by learning [18, 10], or optimization [49] techniques.
+Concurrently, [53] presented a method for jointly performing 2D tracking and 3D lifting, by learning to track with depth supervision while applying an as-rigid-as-possible loss.
+However, their method cannot predict camera poses or identify static parts directly.
+5 Conclusions and limitations
+We presented TracksTo4D, a novel deep-learning framework that directly maps 2D motion tracks from casual videos into their corresponding dynamic structure and camera motion.
+Our approach features a deep learning architecture that considers the symmetries in the problem with designed intrinsic constraints to handle the ill-posed nature of this problem.
+Notably, our network was trained using only raw supervision of 2D point tracks extracted by an off-the-shelf method [18] without any 3D supervision.
+Yet, it implicitly learned to predict camera poses and 3D structures while identifying the dynamic parts.
+During inference, our method demonstrates significantly faster processing times compared to previous methods while achieving comparable accuracy.
+Furthermore, our network exhibits strong generalization capabilities, performing well even on semantic categories that were not present in the training data.
+Limitations and future work.
+While our experiments demonstrated that our network is efficient, accurate, and capable of generalizing to unseen video categories, there are several limitations and future work directions that we would like to address.
+First, Our method cannot handle videos with too rapid motion, and in general, is limited by the accuracy of the tracking method [18].
 We note that any future improvements with point tracking in terms of accuracy and inference time will directly improve our method as well.
 Our method assumes enough motion parallax to constrain the depth values and fails to generate accurate camera poses without it.
-A future and interesting work would be to try combining depth-from-single-image prior as additional inputs to our network for handling cases with minimal motion parallax and for improving accuracies.
-6 Conclusion
-We presented TracksTo4D, a novel deep-learning framework that directly translates 2D motion tracks from casual videos into their corresponding non-rigid structure and motion.
-Our approach features a deep learning architecture that considers the symmetries in the problem with designed intrinsic constraints to handle the ill-posed nature of this problem.
-Notably, our network was trained using only raw supervision of 2D point tracks extracted by an off-the-shelf method [18] without any 3D supervision.
-Yet, it implicitly learned to predict camera poses and non-rigid structures while identifying the non-rigid parts.
-During inference, our method demonstrates faster processing times compared to previous methods while achieving comparable accuracy.
-Furthermore, our network exhibits strong generalization capabilities, performing well even on semantic categories that were not present in the training data.
+A future and interesting work would be to try combining depth-from-single-image prior as additional inputs to our network for handling cases with minimal motion parallax and improving accuracies.
 Acknowledgments
 HM is the Robert J. Shillman Fellow, and is supported by the Israel Science Foundation through a personal grant (ISF 264/23) and an equipment grant (ISF 532/23).
-Appendix 0.A Supplementary Results
-0.A.1 Video Results
+Appendix A Supplementary results
+A.1 Video results
 We provide supplementary video outputs of several cases from our test set.
@@
 Note that the outputs presented in the videos were obtained at inference time, with a single feed-forward prediction, without any optimization or fine-tuning, on unseen test cases.
-0.A.2 Extended Quantitative Evaluation
+A.2 Extended quantitative evaluation
 Per-sequence and mean quantitative comparisons for our 21 pet test videos are presented in Tab.
-5 and Tab.
-6.
-We also show per-sequence fine-tuned results of 2000 iterations starting from our final checkpoint.
-The fine-tuning is done as a post-processing by minimizing the original loss function on the given test video.
-This takes about $10$ minutes per case which is still fast compared to the run time of the previous methods that predicted dynamic depths (see Tab.
-1).
-0.A.3 Generalization Beyond Pet Videos
-We further applied our pretrained network on the MPI Sintel dataset [9] for the evaluation set defined by [64] of $14$ sequences.
-We note that the pet domain that was used for training our network is very different from the Sintel domain which contains fast motion of multiple characters.
-Yet, we can see that due to the strong prior collected by our network, it generalizes to some extent to the domain of Sintel.
-Fig.
-6 shows a qualitative evaluation of our movement level predictions ($\gamma$).
-As can be seen, despite the domain gap (compared to Fig.
-5), our network identifies the non-rigid movement in the videos.
-We further evaluated the camera pose accuracy in Tab.
-7.
-Even though our mean accuracy is not the best among all methods, our method gets the most accurate results on some metrics for several sequences.
-Training on a video dataset in a similar domain to Sintel could possibly improve the generalization of our method for Sintel.
-Unfortunately, the Sintel dataset contains only 23 sequences.
-Figure 6: $\gamma$ Visualization for Sintel.
-We show a visualization of the $\gamma$ outputs of our network.
-In each video sequence, we show the input tracks, where each color visualizes its movement level value, $\gamma$.
-Purple marks rigid points with low $\gamma$ whereas red marks non-rigid points with high $\gamma$.
-Note that our network was only trained on pet videos, and did not see any videos in the domain of Sintel.
-Yet, at inference time with no further optimization, it identifies the non-rigid movement successfully.
-Appendix 0.B Implementation Details
+LABEL:tab::depth_per_seq and Tab.
+LABEL:tab::cam_per_seq.
+Tables with similar structure for the out-of-domain dataset are presented in Tab.
+LABEL:tab::depth_per_seq_nvidia and Tab.
+LABEL:tab::cam_per_seq_nvidia.
+Appendix B Implementation details
 Grid ATE $\downarrow$ RPE Trans $\downarrow$ RPE Rot $\downarrow$ Inference $\downarrow$
@@
 As can be seen, our method can handle a smaller number of points but the accuracy slightly drops with fewer sampled points
-0.B.1 Architecture
-Here we provide a precise definition for our equivariant layer: $F:\mathbb{R}^{N\times P\times d}\rightarrow\mathbb{R}^{N\times P\times d}$.
-The computation of $F$ is split into four steps, as described below:
-$\mathbf{\bar{q}}_{ij}=\bar{W}^{Q}M_{ij},~{}\mathbf{\bar{k}}_{ij}=\bar{W}^{K}M_{ij},~{}\mathbf{\bar{v}}_{ij}=\bar{W}^{V}M_{ij},$ (12a)
-$\bar{M}_{ij}=\sum_{i^{\prime}=1}^{N}\frac{\exp(\mathbf{\bar{q}}_{ij}\cdot\mathbf{\bar{k}}_{i^{\prime}j})}{\sum_{l=1}^{N}\exp(\mathbf{\bar{q}}_{ij}\cdot\mathbf{\bar{k}}_{lj})}\mathbf{\bar{v}}_{i^{\prime}j}$ (12b)
-$\mathbf{q}_{ij}=W^{Q}\bar{M}_{ij},~{}\mathbf{k}_{ij}=W^{K}\bar{M}_{ij},~{}\mathbf{v}_{ij}=W^{V}\bar{M}_{ij},$ (12c)
-$F(M)_{ij}=\sum_{j^{\prime}=1}^{P}\frac{\exp(\mathbf{q}_{ij}\cdot\mathbf{k}_{ij^{\prime}})}{\sum_{l=1}^{P}\exp(\mathbf{q}_{ij}\cdot\mathbf{k}_{il})}\mathbf{v}_{ij^{\prime}}$ (12d)
-Here, $M_{i,j}\in\mathbb{R}^{d}$ are the features associated with the $j$-th point in the $i$-th frame.
-The attention mechanism defined in the first two equations above (with additional temporal positional encoding) replaces the translation equivariant function $L$ applied to the columns of $M$ (Eq.(3)), and the attention in the last two equations above implements the point cloud aggregation (summation) (also in Eq.(3)) applied to the rows of $M$.
-As commonly done we use 16 attention head transformations [46].
-We add standard temporal positional encoding for the transformers with respect to the $N$ (time) dimension as described in [46].
-In our implementation, we added positional encoding only in the first layer.
-Architecture Technical Details
+B.1 Architecture technical details
 For learning high frequencies we map each input coordinate to sinusoidal functions as in [28] with $L=12$.
@@
 We then average over the rows to get per-point features $P_{0}\in\mathbb{R}^{P\times 256}$ and over the columns to get per-frame features $F_{0}\in\mathbb{R}^{N\times 256}$.
-Finally, we map $P_{0}$ to per-point outputs $P_{1}\in\mathbb{R}^{P\times(3K+1)}$ ( K basis points and $\gamma$) with a linear layer, and $F_{0}$ into per-camera outputs $F_{1}\in\mathbb{R}^{N\times(6+3+K)}$ ($6$ for the rotation parameters [67], $3$ for the camera center, and K coefficients) using a conv layer with a kernel size of $31$.
-0.B.2 Other Implementation Details
+Finally, we map $P_{0}$ to per-point outputs $P_{1}\in\mathbb{R}^{P\times(3K+1)}$ ( K basis points and $\gamma$) with a linear layer, and $F_{0}$ into per-camera outputs $F_{1}\in\mathbb{R}^{N\times(6+3+K-1)}$ ($6$ for the rotation parameters [64], $3$ for the camera center, and $K-1$ linear coefficients) using a convolutional layer with a kernel size of $31$.
+B.2 Training details
+In total, we used $733$ cat videos and $753$ dog videos for training.
+We trained our networks for $7000$ and $3500$ epochs for the single-class and multi-class setups respectively.
+Training our method lasts about one week on a single Tesla V100 GPU with 32GB memory.
+We used the Adam optimizer [20] with a learning rate of $10^{-4}$.
+Our method assumes known camera internal parameters which are provided by the dataset and used to normalize the point tracks as a preprocessing step.
+B.3 Other implementation details
 Point tracks sampling
@@
 $20\leq N\leq 50$ and $P=100$.
-When sampling cameras and point tracks of size $N\times P$ from a larger matrix of size $N^{\prime}\times P^{\prime}$ we only take a point track if its starting tracking time is in the range $[t-\frac{N}{2},t+\frac{3N}{2}]$, where t is the first sampled index.
+When sampling cameras and point tracks of size $N\times P\times 3$ from a larger tensor of size $N^{\prime}\times P^{\prime}\times 3$ we only take a point track if its starting tracking time is in the range $[t-\frac{N}{2},t+\frac{3N}{2}]$, where t is the first sampled index.
 At inference time we take all the available point tracks.
 In both, training and inference time, we keep only point tracks that are observed in more than $10$ frames.
+Finetuning details
+For our fine-tuning (FT) in the main paper, we applied per-sequence fine-tunning of 500,100 iterations starting from our final checkpoint, for pets,out-of-domain data respectively.
+The fine-tuning is done as a post-processing by minimizing the original loss function on the given test video.
 Test Set
@@
 They were only used for evaluation and not used by our method at all.
+We verified that this data includes enough dynamic motion, by also including several videos that COLMAP failed to reconstruct without the masks.
+We further verified manually that the camera trajectories look reasonable.
+The out-of-domain dataset contains video sequences with 24 frames, each of resolution of $546\times 288$.
+The GT dynamic masks are provided by the dataset.
+For this dataset, for evaluation only, we define a point track as dynamic if its associated GT mask value is $1$ for at least $15$ frames.
 Bundle Adjustment
-After inference, as optional refinement, we take the output rigid approximation $B_{1}\in\mathbb{R}^{P\times 3}$ and the output camera poses $\{R_{1},\dots,R_{N}\}$, $\{\mathbf{t}_{1},\dots,\mathbf{t}_{N}\}$ and apply Bundle Adjustment (BA).
-We use a 3D world point from $B_{1}$ if its associated $\gamma$ is below $0.008$.
+After inference, as optional refinement, we take the output static approximation $B_{1}\in\mathbb{R}^{P\times 3}$ and the output camera poses $\{R_{1},\dots,R_{N}\}$, $\{\mathbf{t}_{1},\dots,\mathbf{t}_{N}\}$ and apply Bundle Adjustment (BA).
+We use a 3D world point from $B_{1}$ if its associated $\gamma$ is below $0.008$ for the pets dataset and $0.005$ for out-of-domain dataset.
 We optimize reprojection errors of a given observation $M_{i,j}$, only if it is observed, i.e.
@@
 We use the BA implementation provided by [29], which is based on the Ceres package [2].
+Running times
+All inference running times were computed on a machine with NVIDIA RTX A6000 GPU and Intel(R) Core(TM) i7-9800X 3.80GHz CPU.
+Extracting point tracks with [18] took 8.6 and 2.5 seconds for each video on the pet-videos and out-of-domain videos respectively and included in the running time tables as part of our method inference time.
 Training technical details
-In all training setups, we used: $\lambda_{\text{Reprojection}}=50.0$, $\lambda_{\text{Rigid}}=1.0$, $\lambda_{\text{Negative}}=1.0$, $\lambda_{\text{Sparse}}=0.001$.
+In all training setups, we used: $\lambda_{\text{Reprojection}}=50.0$, $\lambda_{\text{Static}}=1.0$, $\lambda_{\text{Negative}}=1.0$, $\lambda_{\text{Sparse}}=0.001$.
 At the beginning of the training, we pre-train the camera poses to be located behind and facing the origin.
@@
 many points have negative depths, which may result in bad convergence.
-More specifically, the pre-train loss is:
-$\mathcal{L}_{\text{Pretrain}}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{100}\left\lVert\mathbf{t}_{i}-[0,0,-15]^{T}\right\rVert^{2}+\left\lVert R_{i}-I\right\rVert_{F}^{2}$
+More specifically, the pre-train loss is: $\mathcal{L}_{\text{Pretrain}}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{100}\left\lVert\mathbf{t}_{i}-[0,0,-15]^{T}\right\rVert^{2}+\left\lVert R_{i}-I\right\rVert_{F}^{2}$.
 The pretrain runs until convergence ($\mathcal{L}_{\text{Pretrain}}<10^{-4})$.
@@
 Until epoch $50$ we sample sequences of $N$ in the range $[20,22]$, and then we increase the range to $[20,50]$.
-Table 5: Depth Accuracy.
+Table 5: Depth accuracy, pet test set.
 We show a comparison to previous methods on the predicted depth for the point trajectories compared to their GT depths.
@@
 $\delta<1.25^{3}\uparrow$ Dyn 1.00 1.00 1.00 1.00 1.00 1.00 1.00
-All 0.89 0.91 0.97 1.00 0.99 1.00 0.99
+All 0.89 0.91 0.97 1.00 1.00 1.00 1.00
 […]
 All 0.19 0.18 0.09 0.09 0.07 0.09 0.07
-$\delta<1.25\uparrow$ Dyn 0.45 0.75 0.81 0.81 0.85 0.90 0.87
-All 0.68 0.76 0.92 0.90 0.91 0.94 0.92
+$\delta<1.25\uparrow$ Dyn 0.45 0.75 0.81 0.81 0.86 0.90 0.87
+All 0.68 0.76 0.92 0.90 0.92 0.94 0.92
 $\delta<1.25^{2}\uparrow$ Dyn 0.85 0.99 0.89 1.00 1.00 0.99 1.00
@@
 All 0.22 0.24 0.06 0.09 0.06 0.08 0.06
-$\delta<1.25\uparrow$ Dyn 0.10 0.53 0.97 0.95 0.98 0.97 0.98
-All 0.76 0.74 0.97 0.91 0.97 0.92 0.96
+$\delta<1.25\uparrow$ Dyn 0.10 0.53 0.97 0.95 0.99 0.97 0.98
+All 0.76 0.74 0.97 0.91 0.97 0.92 0.97
 $\delta<1.25^{2}\uparrow$ Dyn 0.27 0.99 1.00 1.00 1.00 1.00 1.00
@@
 All 0.17 0.27 0.09 0.12 0.11 0.10 0.11
-$\delta<1.25\uparrow$ Dyn 0.44 0.98 1.00 0.54 0.75 0.65 0.76
+$\delta<1.25\uparrow$ Dyn 0.44 0.98 1.00 0.54 0.75 0.65 0.77
 All 0.80 0.65 0.96 0.84 0.92 0.89 0.92
@@
 […]
-All 0.12 3.75E+05 0.03 0.08 0.04 0.06 0.04
-$\delta<1.25\uparrow$ Dyn 0.97 0.98 1.00 0.99 0.98 0.99 0.99
-All 0.86 0.86 0.98 0.91 0.98 0.96 0.98
+All 0.12 3.75E+05 0.03 0.08 0.05 0.06 0.04
+$\delta<1.25\uparrow$ Dyn 0.97 0.98 1.00 0.99 0.98 0.99 0.98
+All 0.86 0.86 0.98 0.91 0.97 0.96 0.98
 $\delta<1.25^{2}\uparrow$ Dyn 1.00 1.00 1.00 1.00 1.00 1.00 1.00
@@
 […]
-All 0.14 0.15 0.05 0.04 0.06 0.05 0.06
+All 0.14 0.15 0.05 0.04 0.05 0.05 0.05
 $\delta<1.25\uparrow$ Dyn 0.47 0.94 0.98 0.99 0.99 0.99 0.99
@@
 All 0.22 0.17 0.06 0.10 0.06 0.09 0.06
-$\delta<1.25\uparrow$ Dyn 0.45 0.95 0.92 0.84 0.89 0.90 0.91
-All 0.65 0.80 0.97 0.86 0.96 0.91 0.96
-$\delta<1.25^{2}\uparrow$ Dyn 0.68 1.00 0.99 0.98 0.99 0.98 0.99
+$\delta<1.25\uparrow$ Dyn 0.45 0.95 0.92 0.84 0.91 0.90 0.91
+All 0.65 0.80 0.97 0.86 0.96 0.91 0.97
+$\delta<1.25^{2}\uparrow$ Dyn 0.68 1.00 0.99 0.98 0.98 0.98 0.98
 All 0.86 0.92 0.99 0.98 0.99 0.98 0.99
-$\delta<1.25^{3}\uparrow$ Dyn 0.93 1.00 1.00 1.00 1.00 1.00 1.00
+$\delta<1.25^{3}\uparrow$ Dyn 0.93 1.00 1.00 1.00 0.99 1.00 0.99
 All 0.97 0.98 1.00 0.99 1.00 0.99 1.00
@@
 All 0.33 0.20 0.09 0.21 0.09 0.22 0.09
-$\delta<1.25\uparrow$ Dyn 0.05 0.59 0.80 0.46 0.60 0.45 0.59
-All 0.65 0.74 0.92 0.55 0.88 0.52 0.88
+$\delta<1.25\uparrow$ Dyn 0.05 0.59 0.80 0.46 0.63 0.45 0.60
+All 0.65 0.74 0.92 0.55 0.88 0.52 0.87
 $\delta<1.25^{2}\uparrow$ Dyn 0.28 0.85 0.95 0.91 1.00 0.85 1.00
-All 0.78 0.91 0.98 0.89 0.99 0.86 0.99
+All 0.78 0.91 0.98 0.89 0.99 0.86 0.98
 $\delta<1.25^{3}\uparrow$ Dyn 0.68 1.00 1.00 1.00 1.00 1.00 1.00
-All 0.90 0.98 1.00 0.97 1.00 0.97 1.00
+All 0.90 0.98 1.00 0.97 1.00 0.97 0.99
 […]
@@
 $\delta<1.25^{2}\uparrow$ Dyn 0.86 1.00 0.99 0.96 0.98 0.97 0.98
-All 0.94 0.78 0.99 0.99 1.00 0.99 0.99
+All 0.94 0.78 0.99 0.99 0.99 0.99 0.99
 $\delta<1.25^{3}\uparrow$ Dyn 0.94 1.00 1.00 0.99 1.00 0.99 1.00
@@
 […]
-All 0.16 0.13 0.05 0.08 0.08 0.08 0.08
+All 0.16 0.13 0.05 0.08 0.06 0.08 0.07
 $\delta<1.25\uparrow$ Dyn 0.19 0.79 0.98 0.98 0.97 0.98 0.97
@@
 $\delta<1.25^{2}\uparrow$ Dyn 0.97 1.00 1.00 1.00 1.00 1.00 1.00
-All 0.98 0.98 1.00 0.99 0.97 0.99 0.97
+All 0.98 0.98 1.00 0.99 0.98 0.99 0.98
 $\delta<1.25^{3}\uparrow$ Dyn 1.00 1.00 1.00 1.00 1.00 1.00 1.00
-All 1.00 1.00 1.00 1.00 0.98 1.00 0.98
+All 1.00 1.00 1.00 1.00 1.00 1.00 1.00
 […]
-All 3.31E+08 0.35 0.06 0.10 0.08 0.09 0.07
-$\delta<1.25\uparrow$ Dyn 0.50 0.50 0.93 0.82 0.95 0.85 0.96
-All 0.67 0.69 0.97 0.91 0.97 0.93 0.97
+All 3.31E+08 0.35 0.06 0.10 0.07 0.09 0.07
+$\delta<1.25\uparrow$ Dyn 0.50 0.50 0.93 0.82 0.92 0.85 0.95
+All 0.67 0.69 0.97 0.91 0.96 0.93 0.97
 $\delta<1.25^{2}\uparrow$ Dyn 0.81 0.84 1.00 1.00 1.00 1.00 1.00
@@
 […]
-All 0.33 0.39 0.09 0.15 0.11 0.18 0.10
-$\delta<1.25\uparrow$ Dyn 0.01 0.77 0.95 0.62 0.90 0.79 0.88
-All 0.65 0.65 0.97 0.80 0.93 0.70 0.93
+All 0.33 0.39 0.09 0.15 0.09 0.18 0.09
+$\delta<1.25\uparrow$ Dyn 0.01 0.77 0.95 0.62 0.96 0.79 0.97
+All 0.65 0.65 0.97 0.80 0.94 0.70 0.95
 $\delta<1.25^{2}\uparrow$ Dyn 0.12 1.00 0.99 0.93 0.99 0.96 0.99
-All 0.81 0.83 0.99 0.97 0.98 0.92 0.98
-$\delta<1.25^{3}\uparrow$ Dyn 0.46 1.00 0.99 1.00 0.99 1.00 0.99
-All 0.90 0.89 0.99 0.99 0.98 0.99 0.99
+All 0.81 0.83 0.99 0.97 0.99 0.92 0.99
+$\delta<1.25^{3}\uparrow$ Dyn 0.46 1.00 0.99 1.00 1.00 1.00 1.00
+All 0.90 0.89 0.99 0.99 0.99 0.99 0.99
 […]
 All 0.21 0.21 0.05 0.06 0.06 0.06 0.06
-$\delta<1.25\uparrow$ Dyn 0.42 0.92 0.98 0.81 0.89 0.80 0.90
+$\delta<1.25\uparrow$ Dyn 0.42 0.92 0.98 0.81 0.88 0.80 0.90
 All 0.66 0.79 0.99 0.96 0.97 0.96 0.97
@@
 […]
-All 0.23 0.15 0.08 0.07 0.07 0.07 0.06
-$\delta<1.25\uparrow$ Dyn 0.40 0.88 0.76 0.93 0.87 0.91 0.88
+All 0.23 0.15 0.08 0.07 0.06 0.07 0.07
+$\delta<1.25\uparrow$ Dyn 0.40 0.88 0.76 0.93 0.89 0.91 0.88
 All 0.60 0.81 0.89 0.94 0.94 0.93 0.94
@@
 All 0.87 0.94 0.98 0.98 0.99 0.99 0.99
-$\delta<1.25^{3}\uparrow$ Dyn 0.95 1.00 1.00 0.99 1.00 0.99 1.00
+$\delta<1.25^{3}\uparrow$ Dyn 0.95 1.00 1.00 0.99 0.99 0.99 0.99
 All 0.95 0.98 1.00 0.99 1.00 1.00 1.00
@@
 All 0.14 0.32 0.05 0.10 0.09 0.08 0.09
-$\delta<1.25\uparrow$ Dyn 0.46 0.60 0.93 0.87 0.95 0.88 0.95
+$\delta<1.25\uparrow$ Dyn 0.46 0.60 0.93 0.87 0.96 0.88 0.96
 All 0.86 0.65 0.98 0.91 0.96 0.95 0.96
@@
 All 2.42E+08 0.45 0.04 0.05 0.05 0.05 0.05
-$\delta<1.25\uparrow$ Dyn 0.34 0.49 0.61 0.44 0.38 0.49 0.34
+$\delta<1.25\uparrow$ Dyn 0.34 0.49 0.61 0.44 0.37 0.49 0.34
 All 0.50 0.42 0.97 0.95 0.95 0.95 0.95
@@
 All 3.63E+07 6.16E+04 0.06 0.08 0.06 0.08 0.06
-$\delta<1.25\uparrow$ Dyn 0.43 0.78 0.93 0.85 0.90 0.88 0.89
+$\delta<1.25\uparrow$ Dyn 0.43 0.78 0.93 0.85 0.90 0.88 0.90
 All 0.72 0.71 0.97 0.91 0.96 0.92 0.96
@@
 $\delta<1.25^{3}\uparrow$ Dyn 0.92 1.00 1.00 1.00 1.00 1.00 1.00
-All 0.96 0.93 1.00 1.00 0.99 1.00 0.99
-Table 6: Camera Poses Accuracy.
+All 0.96 0.93 1.00 1.00 1.00 1.00 1.00
+[…]
+All 9.39E+07 2.04E+05 0.02 0.04 0.02 0.04 0.02
+$\delta<1.25\uparrow$ Dyn 0.29 0.17 0.10 0.18 0.15 0.16 0.16
+All 0.10 0.11 0.03 0.10 0.03 0.11 0.03
+$\delta<1.25^{2}\uparrow$ Dyn 0.26 0.05 0.02 0.03 0.01 0.03 0.01
+All 0.07 0.07 0.01 0.02 0.01 0.03 0.01
+$\delta<1.25^{3}\uparrow$ Dyn 0.15 0.00 0.00 0.00 0.00 0.00 0.00
+All 0.05 0.05 0.01 0.01 0.00 0.01 0.00
+Table 6: Camera poses accuracy for pets.
 We show a comparison to previous methods on the predicted camera poses.
@@
 Ours (C)+BA: Our inference time outputs, followed by a short Bundle Adjustment (total inference time of 0.4 seconds) for a model that was trained only on cats.
-Ours (C) FT: The outputs of the model that was trained only on cats after fine-tuning our losses for each specific video (total running time of about 10 minutes).
+Ours (C) FT: The outputs of the model that was trained only on cats after fine-tuning our losses for each specific video (total running time of about 5 minutes).
 As can be seen, after BA, our results are the most accurate compared to the other methods, and fine-tuning improves our accuracy even more.
 DROID-SLAM[44] ParticleSfM[61] RCVD[22] CasualSAM[59] Ours Ours Ours
-(C) (C)+BA (Csho)+FT
-Seq0 ATE(mm) 3.71 6.10 64.67 5.36 5.60 5.43 4.40
-RPE T.(mm) 3.05 3.22 26.92 3.13 3.63 3.04 2.91
-RPE R.(deg.) 0.14 0.18 2.53 0.16 0.22 0.15 0.13
-Seq1 ATE(mm) 1.91 3.83 38.44 10.28 13.21 4.32 1.71
-RPE T.(mm) 1.32 1.49 25.23 3.00 2.65 1.72 1.29
+(C) (C)+BA (C)+FT
+Seq0 ATE(mm) 3.71 6.10 64.67 5.36 5.60 5.43 4.42
+RPE T.(mm) 3.05 3.22 26.92 3.13 3.63 3.04 2.89
+RPE R.(deg.) 0.14 0.18 2.53 0.16 0.22 0.15 0.14
+Seq1 ATE(mm) 1.91 3.83 38.44 10.28 13.21 4.32 1.76
+RPE T.(mm) 1.32 1.49 25.23 3.00 2.65 1.72 1.32
 RPE R.(deg.) 0.18 0.23 2.43 0.67 0.29 0.19 0.16
-Seq2 ATE(mm) 3.13 4.68 39.46 2.13 4.57 2.78 2.52
-RPE T.(mm) 3.62 5.82 27.27 1.97 3.24 2.62 2.55
+Seq2 ATE(mm) 3.13 4.68 39.46 2.13 4.57 2.78 2.53
+RPE T.(mm) 3.62 5.82 27.27 1.97 3.24 2.62 2.49
 RPE R.(deg.) 0.08 0.21 1.89 0.06 0.12 0.09 0.09
-Seq3 ATE(mm) 5.13 2.26 29.55 2.49 7.83 2.66 2.87
-RPE T.(mm) 5.76 2.00 24.18 2.31 3.13 2.27 2.38
+Seq3 ATE(mm) 5.13 2.26 29.55 2.49 7.83 2.66 2.92
+RPE T.(mm) 5.76 2.00 24.18 2.31 3.13 2.27 2.30
 RPE R.(deg.) 0.17 0.07 2.38 0.08 0.13 0.07 0.07
-Seq4 ATE(mm) 2.59 4.38 56.16 2.65 4.21 2.38 2.24
-RPE T.(mm) 2.05 2.07 21.36 1.74 2.40 2.05 1.95
-RPE R.(deg.) 0.11 0.11 1.05 0.09 0.15 0.11 0.10
-Seq5 ATE(mm) 1.07 0.83 14.22 0.53 1.44 0.79 0.74
+Seq4 ATE(mm) 2.59 4.38 56.16 2.65 4.21 2.38 2.28
+RPE T.(mm) 2.05 2.07 21.36 1.74 2.40 2.05 2.04
+RPE R.(deg.) 0.11 0.11 1.05 0.09 0.15 0.11 0.11
+Seq5 ATE(mm) 1.07 0.83 14.22 0.53 1.44 0.79 0.75
 RPE T.(mm) 1.02 0.74 6.51 0.52 0.88 0.68 0.68
 RPE R.(deg.) 0.09 0.07 1.63 0.05 0.12 0.07 0.06
-Seq6 ATE(mm) 26.08 31.07 48.31 26.12 27.10 28.28 29.80
-RPE T.(mm) 10.57 11.01 24.21 10.31 10.44 10.92 11.02
+Seq6 ATE(mm) 26.08 31.07 48.31 26.12 27.10 28.28 28.98
+RPE T.(mm) 10.57 11.01 24.21 10.31 10.44 10.92 10.82
 RPE R.(deg.) 0.66 0.67 4.03 0.62 0.69 0.69 0.68
-Seq7 ATE(mm) 2.25 28.48 47.25 1.78 4.53 2.39 2.11
-RPE T.(mm) 2.34 6.13 23.81 1.72 2.32 1.98 1.95
-RPE R.(deg.) 0.15 0.59 2.29 0.11 0.20 0.16 0.15
-Seq8 ATE(mm) 1.23 38.79 44.06 2.00 4.78 0.84 0.96
-RPE T.(mm) 1.07 50.57 25.46 1.24 1.54 0.71 0.74
+Seq7 ATE(mm) 2.25 28.48 47.25 1.78 4.53 2.39 2.25
+RPE T.(mm) 2.34 6.13 23.81 1.72 2.32 1.98 1.99
+RPE R.(deg.) 0.15 0.59 2.29 0.11 0.20 0.16 0.16
+Seq8 ATE(mm) 1.23 38.79 44.06 2.00 4.78 0.84 0.89
+RPE T.(mm) 1.07 50.57 25.46 1.24 1.54 0.71 0.70
 RPE R.(deg.) 0.07 4.78 1.73 0.09 0.13 0.06 0.05
-Seq9 ATE(mm) 21.74 18.52 43.45 34.93 24.15 3.92 3.28
-RPE T.(mm) 9.04 7.96 21.35 21.06 6.89 2.77 2.01
+Seq9 ATE(mm) 21.74 18.52 43.45 34.93 24.15 3.92 3.22
+RPE T.(mm) 9.04 7.96 21.35 21.06 6.89 2.77 2.08
 RPE R.(deg.) 0.62 0.47 3.47 0.71 0.37 0.12 0.10
-Seq10 ATE(mm) 1.47 1.75 22.49 1.40 3.03 1.42 1.48
-RPE T.(mm) 1.84 1.15 24.22 1.11 1.87 1.25 1.24
+Seq10 ATE(mm) 1.47 1.75 22.49 1.40 3.03 1.42 1.46
+RPE T.(mm) 1.84 1.15 24.22 1.11 1.87 1.25 1.22
 RPE R.(deg.) 0.22 0.12 2.18 0.10 0.21 0.12 0.12
-Seq11 ATE(mm) 1.71 3.60 19.10 1.32 3.12 1.66 1.50
-RPE T.(mm) 1.65 1.54 16.34 1.21 3.12 1.75 1.60
+Seq11 ATE(mm) 1.71 3.60 19.10 1.32 3.12 1.66 1.49
+RPE T.(mm) 1.65 1.54 16.34 1.21 3.12 1.75 1.55
 RPE R.(deg.) 0.08 0.08 1.80 0.06 0.12 0.08 0.07
-Seq12 ATE(mm) 1.70 3.64 20.82 2.51 6.23 3.61 2.58
-RPE T.(mm) 2.24 2.63 18.50 2.74 4.04 3.02 2.34
+Seq12 ATE(mm) 1.70 3.64 20.82 2.51 6.23 3.61 2.54
+RPE T.(mm) 2.24 2.63 18.50 2.74 4.04 3.02 2.84
 RPE R.(deg.) 0.09 0.12 2.03 0.12 0.23 0.15 0.14
-Seq13 ATE(mm) 1.23 2.38 33.49 1.49 4.10 1.17 1.27
-RPE T.(mm) 1.19 1.40 17.33 1.00 1.72 1.12 1.01
-RPE R.(deg.) 0.13 0.14 2.12 0.12 0.19 0.12 0.11
-Seq14 ATE(mm) 5.42 5.15 1.05E+02 24.95 6.09 5.06 4.90
-RPE T.(mm) 3.40 3.38 69.36 9.45 4.07 3.57 3.55
+Seq13 ATE(mm) 1.23 2.38 33.49 1.49 4.10 1.17 1.28
+RPE T.(mm) 1.19 1.40 17.33 1.00 1.72 1.12 1.09
+RPE R.(deg.) 0.13 0.14 2.12 0.12 0.19 0.12 0.12
+Seq14 ATE(mm) 5.42 5.15 1.05E+02 24.95 6.09 5.06 4.93
+RPE T.(mm) 3.40 3.38 69.36 9.45 4.07 3.57 3.41
 RPE R.(deg.) 0.18 0.19 3.81 0.65 0.26 0.20 0.19
-Seq15 ATE(mm) 7.69 61.06 36.70 7.40 36.51 4.98 5.50
-RPE T.(mm) 7.95 17.57 28.18 6.93 10.19 5.86 5.74
+Seq15 ATE(mm) 7.69 61.06 36.70 7.40 36.51 4.98 5.41
+RPE T.(mm) 7.95 17.57 28.18 6.93 10.19 5.86 5.72
 RPE R.(deg.) 0.22 0.58 2.85 0.19 0.30 0.16 0.15
 Seq16 ATE(mm) 5.04 5.06 36.42 3.69 4.52 4.11 3.81
-RPE T.(mm) 4.53 4.54 20.11 4.02 4.47 4.25 4.25
-RPE R.(deg.) 0.28 0.29 2.94 0.25 0.31 0.28 0.28
-Seq17 ATE(mm) 1.12 34.07 77.91 2.08 6.52 2.51 2.89
-RPE T.(mm) 1.15 12.30 39.64 1.18 2.14 1.50 1.47
+RPE T.(mm) 4.53 4.54 20.11 4.02 4.47 4.25 4.15
+RPE R.(deg.) 0.28 0.29 2.94 0.25 0.31 0.28 0.27
+Seq17 ATE(mm) 1.12 34.07 77.91 2.08 6.52 2.51 2.86
+RPE T.(mm) 1.15 12.30 39.64 1.18 2.14 1.50 1.55
 RPE R.(deg.) 0.10 1.11 2.20 0.08 0.17 0.13 0.13
-Seq18 ATE(mm) 2.98 6.25 36.54 4.91 7.73 4.13 3.83
-RPE T.(mm) 4.05 5.21 24.78 3.89 4.34 3.62 3.80
+Seq18 ATE(mm) 2.98 6.25 36.54 4.91 7.73 4.13 3.86
+RPE T.(mm) 4.05 5.21 24.78 3.89 4.34 3.62 3.65
 RPE R.(deg.) 0.23 0.29 1.67 0.21 0.24 0.21 0.21
-Seq19 ATE(mm) 1.45 2.23 42.95 1.82 8.51 2.79 2.39
-RPE T.(mm) 1.65 1.72 29.12 1.44 3.23 2.21 1.97
-RPE R.(deg.) 0.11 0.11 2.09 0.09 0.25 0.16 0.14
-Seq20 ATE(mm) 8.13 4.37 66.03 4.95 4.30 3.43 4.03
-RPE T.(mm) 6.20 3.57 27.34 3.00 3.30 3.05 3.08
+Seq19 ATE(mm) 1.45 2.23 42.95 1.82 8.51 2.79 2.36
+RPE T.(mm) 1.65 1.72 29.12 1.44 3.23 2.21 2.01
+RPE R.(deg.) 0.11 0.11 2.09 0.09 0.25 0.16 0.15
+Seq20 ATE(mm) 8.13 4.37 66.03 4.95 4.30 3.43 4.06
+RPE T.(mm) 6.20 3.57 27.34 3.00 3.30 3.05 3.06
 RPE R.(deg.) 0.36 0.20 1.35 0.18 0.20 0.18 0.18
-Mean ATE(mm) 5.08 12.79 43.95 6.90 8.96 4.22 4.04
+Mean ATE(mm) 5.08 12.79 43.95 6.90 8.96 4.22 4.00
 RPE T.(mm) 3.60 6.95 25.77 3.95 3.79 2.86 2.74
 RPE R.(deg.) 0.20 0.51 2.31 0.22 0.23 0.17 0.16
-Table 7: Camera Poses Accuracy on Sintel.
+STD ATE(mm) 6.63 16.32 21.22 9.55 9.06 5.68 5.87
+RPE T.(mm) 2.80 10.88 11.80 4.74 2.52 2.22 2.22
+RPE R.(deg.) 0.16 1.01 0.76 0.22 0.13 0.13 0.13
+Table 7: Depth accuracy for out-of-domain data [58].
+We show a comparison to previous methods on the predicted depth for the point trajectories compared to their GT depths.
+We compare 2 ways of running our method.
+Ours (C&D): Our inference time outputs for a model that was trained on cats and dogs.
+Ours (C&D) FT: The outputs of our model trained on cats and dogs, after fine-tuning our losses for each specific video.
+As can be seen, fine-tuning improves our accuracy even more.
+RCVD [22] MiDaS[4] CasualSAM[59] Ours (C&D) Ours (C&D) FT
+[…]
+All 0.14 0.34 0.02 0.06 0.01
+$\delta<1.25\uparrow$ Dyn 0.42 0.89 0.98 0.98 0.98
+All 0.62 0.72 0.99 0.99 0.99
+$\delta<1.25^{2}\uparrow$ Dyn 1.00 0.98 0.99 0.99 0.99
+All 1.00 0.81 1.00 1.00 1.00
+$\delta<1.25^{3}\uparrow$ Dyn 1.00 0.99 1.00 1.00 1.00
+All 1.00 0.88 1.00 1.00 1.00
+[…]
+All 0.14 4.87E+05 0.01 0.06 0.01
+$\delta<1.25\uparrow$ Dyn 0.97 0.95 1.00 0.99 1.00
+All 0.83 0.76 1.00 0.97 1.00
+$\delta<1.25^{2}\uparrow$ Dyn 1.00 0.99 1.00 1.00 1.00
+All 1.00 0.86 1.00 1.00 1.00
+$\delta<1.25^{3}\uparrow$ Dyn 1.00 0.99 1.00 1.00 1.00
+All 1.00 0.93 1.00 1.00 1.00
+[…]
+All 0.05 4.75E+04 0.01 0.06 0.01
+$\delta<1.25\uparrow$ Dyn 0.94 0.03 0.99 0.98 0.98
+All 0.98 0.67 1.00 1.00 1.00
+$\delta<1.25^{2}\uparrow$ Dyn 1.00 0.04 1.00 0.98 1.00
+All 1.00 0.82 1.00 1.00 1.00
+$\delta<1.25^{3}\uparrow$ Dyn 1.00 0.18 1.00 1.00 1.00
+All 1.00 0.86 1.00 1.00 1.00
+[…]
+All 0.12 0.59 0.02 0.05 0.04
+$\delta<1.25\uparrow$ Dyn 0.77 0.07 0.99 0.96 0.96
+All 0.86 0.22 0.99 0.97 0.97
+$\delta<1.25^{2}\uparrow$ Dyn 1.00 0.30 1.00 1.00 0.99
+All 1.00 0.38 1.00 1.00 1.00
+$\delta<1.25^{3}\uparrow$ Dyn 1.00 0.75 1.00 1.00 1.00
+All 1.00 0.64 1.00 1.00 1.00
+[…]
+All 0.30 7.67E+03 0.07 0.15 0.08
+$\delta<1.25\uparrow$ Dyn 0.36 0.49 0.93 0.64 0.89
+All 0.44 0.59 0.96 0.78 0.94
+$\delta<1.25^{2}\uparrow$ Dyn 0.61 0.72 0.99 0.98 0.94
+All 0.71 0.78 0.98 0.91 0.97
+$\delta<1.25^{3}\uparrow$ Dyn 0.67 0.83 1.00 0.98 0.94
+All 0.82 0.87 0.99 0.98 0.98
+[…]
+All 0.10 1.09 0.02 0.05 0.04
+$\delta<1.25\uparrow$ Dyn 0.89 0.59 0.76 0.93 0.93
+All 0.92 0.29 0.99 1.00 0.99
+$\delta<1.25^{2}\uparrow$ Dyn 1.00 0.95 0.97 0.97 0.97
+All 1.00 0.42 1.00 1.00 1.00
+$\delta<1.25^{3}\uparrow$ Dyn 1.00 0.97 1.00 1.00 0.97
+All 1.00 0.53 1.00 1.00 1.00
+[…]
+All 2.06E+06 0.22 0.03 0.14 0.05
+$\delta<1.25\uparrow$ Dyn 0.26 0.81 0.99 0.94 1.00
+All 0.50 0.71 1.00 0.79 1.00
+$\delta<1.25^{2}\uparrow$ Dyn 0.99 0.99 1.00 1.00 1.00
+All 0.88 0.94 1.00 1.00 1.00
+$\delta<1.25^{3}\uparrow$ Dyn 1.00 1.00 1.00 1.00 1.00
+All 1.00 0.99 1.00 1.00 1.00
+[…]
+All 0.10 1.62E+06 0.02 0.05 0.02
+$\delta<1.25\uparrow$ Dyn 0.91 0.87 1.00 1.00 1.00
+All 0.84 0.68 1.00 0.99 1.00
+$\delta<1.25^{2}\uparrow$ Dyn 1.00 0.91 1.00 1.00 1.00
+All 1.00 0.71 1.00 1.00 1.00
+$\delta<1.25^{3}\uparrow$ Dyn 1.00 0.92 1.00 1.00 1.00
+All 1.00 0.72 1.00 1.00 1.00
+[…]
+All 2.58E+05 2.70E+05 0.03 0.08 0.03
+$\delta<1.25\uparrow$ Dyn 0.69 0.59 0.95 0.93 0.97
+All 0.75 0.58 0.99 0.94 0.99
+$\delta<1.25^{2}\uparrow$ Dyn 0.95 0.73 0.99 0.99 0.99
+All 0.95 0.72 1.00 0.99 1.00
+$\delta<1.25^{3}\uparrow$ Dyn 0.96 0.83 1.00 1.00 0.99
+All 0.98 0.80 1.00 1.00 1.00
+[…]
+All 7.28E+05 5.69E+05 0.02 0.04 0.02
+$\delta<1.25\uparrow$ Dyn 0.29 0.37 0.08 0.12 0.04
+All 0.20 0.21 0.01 0.10 0.02
+$\delta<1.25^{2}\uparrow$ Dyn 0.14 0.37 0.01 0.01 0.02
+All 0.10 0.21 0.01 0.03 0.01
+$\delta<1.25^{3}\uparrow$ Dyn 0.12 0.28 0.00 0.01 0.02
+All 0.06 0.16 0.00 0.01 0.01
+Table 8: Camera poses accuracy for out-of-domain data [58].
 We show a comparison to previous methods on the predicted camera poses.
-We compare 4 ways of running our method.
-Ours (C), Ours (C&D): Our inference time outputs for a model that was trained only on cats or on cats and dogs respectively.
-Ours (C)+BA, Ours (C&D)+BA: The outputs of the model trained only on cats or on cats and dogs respectively, after further applying short BA refinement.
-The numbers for [21],[43] and [64] were obtained from [64].
-RCVD[21] DROID- Particle CasualSAM[62] Ours Ours Ours Ours
-SLAM[43] SfM[64] (C) (C)+BA (C&D) (C&D)+BA
-alley_2 ATE(mm) 0.026 0.057 0.002 0.0008 0.043 0.027 0.055 0.046
-RPE T.(mm) 0.056 0.035 0.0007 0.0004 0.033 0.040 0.034 0.041
-RPE R.(deg.) 0.821 1.047 0.009 0.008 0.468 0.507 0.511 0.531
-ambush_4 ATE(mm) 0.171 0.104 0.068 0.011 0.050 0.021 0.055 0.022
-RPE T.(mm) 0.048 0.035 0.027 0.007 0.026 0.026 0.026 0.026
-RPE R.(deg.) 3.025 1.385 0.473 0.079 0.398 0.319 0.382 0.324
-ambush_5 ATE(mm) 0.230 0.112 0.002 0.008 0.057 0.010 0.035 0.009
-RPE T.(mm) 0.046 0.029 0.001 0.003 0.018 0.008 0.015 0.006
-RPE R.(deg.) 4.105 1.580 0.055 0.084 0.430 0.228 0.338 0.149
-ambush_6 ATE(mm) 0.199 0.289 0.269 0.042 0.241 0.223 0.191 0.141
-RPE T.(mm) 0.112 0.078 0.081 0.024 0.088 0.129 0.072 0.078
-RPE R.(deg.) 4.147 4.596 0.951 0.482 1.928 3.364 1.602 1.849
-cave_2 ATE(mm) 0.596 0.351 0.961 0.967 0.217 0.217 0.460 0.460
-RPE T.(mm) 0.171 0.172 0.142 0.170 0.094 0.094 0.113 0.113
-RPE R.(deg.) 7.508 5.489 3.678 2.674 1.781 1.781 2.634 2.634
-cave_4 ATE(mm) 0.179 0.155 0.068 0.009 0.055 0.006 0.115 0.025
-RPE T.(mm) 0.087 0.035 0.012 0.003 0.024 0.006 0.028 0.021
-RPE R.(deg.) 2.040 2.710 0.409 0.052 0.244 0.044 0.294 0.171
-market_2 ATE(mm) 0.032 0.011 0.003 0.001 0.027 0.002 0.033 0.003
-RPE T.(mm) 0.018 0.006 0.010 0.0006 0.004 0.002 0.007 0.002
-RPE R.(deg.) 0.141 0.036 0.041 0.010 0.107 0.008 0.091 0.008
-market_5 ATE(mm) 1.213 0.912 0.012 0.005 0.891 1.187 0.859 0.525
-RPE T.(mm) 0.762 0.293 0.006 0.002 0.212 0.557 0.209 0.395
-RPE R.(deg.) 1.863 3.334 0.034 0.016 1.252 6.946 1.489 3.285
-market_6 ATE(mm) 0.248 0.057 0.018 0.008 0.187 0.163 0.220 0.152
-RPE T.(mm) 0.214 0.037 0.007 0.002 0.103 0.097 0.095 0.086
-RPE R.(deg.) 0.817 1.296 0.120 0.019 1.003 0.850 0.834 0.695
-shaman_3 ATE(mm) 0.054 0.001 0.0005 0.002 0.002 0.002 0.003 0.002
-RPE T.(mm) 0.023 0.002 0.0004 0.0004 0.0007 0.0004 0.001 0.0004
-RPE R.(deg.) 0.718 0.199 0.072 0.036 0.060 0.042 0.087 0.039
-sleeping_1 ATE(mm) 0.029 0.011 0.008 0.008 0.002 0.002 0.002 0.002
-RPE T.(mm) 0.019 0.006 0.001 0.001 0.001 0.0006 0.001 0.0006
-RPE R.(deg.) 0.668 0.479 0.042 0.040 0.070 0.040 0.096 0.040
-sleeping_2 ATE(mm) 0.043 0.005 0.0007 0.0002 0.005 0.0006 0.005 0.0006
-RPE T.(mm) 0.049 0.018 0.0002 0.0002 0.003 0.0007 0.003 0.0007
-RPE R.(deg.) 0.446 0.139 0.006 0.008 0.108 0.018 0.088 0.018
-temple_2 ATE(mm) 1.245 0.073 0.011 0.086 0.189 0.011 0.217 0.009
-RPE T.(mm) 0.394 0.348 0.002 0.008 0.029 0.004 0.032 0.003
-RPE R.(deg.) 1.318 1.298 0.019 0.131 0.287 0.019 0.340 0.018
-temple_3 ATE(mm) 0.769 0.310 0.381 0.587 0.350 0.510 0.329 0.473
-RPE T.(mm) 0.161 0.093 0.149 0.136 0.114 0.190 0.114 0.138
-RPE R.(deg.) 20.592 3.230 1.583 0.109 1.672 2.110 1.458 1.176
-Mean ATE(mm) 0.360 0.175 0.129 0.124 0.165 0.170 0.184 0.133
-RPE T.(mm) 0.154 0.085 0.031 0.025 0.054 0.082 0.054 0.065
-RPE R.(deg.) 3.444 1.916 0.535 0.268 0.701 1.163 0.732 0.781
+We compare 3 ways of running our method.
+Ours (C): Our inference time outputs.
+Ours (C)+BA: Our inference time outputs, followed by a short Bundle Adjustment for a model that was trained only on cats.
+Ours (C) FT: The outputs of the model that was trained only on cats after fine-tuning our losses for each specific video
+DROID-SLAM[44] ParticleSfM[61] RCVD[22] CasualSAM[59] Ours Ours Ours
+(C) (C)+BA (C)+FT
+Balloon1 ATE(mm) 2.87 5.81 1.7E+02 5.57 21.47 4.17 4.14
+RPE T.(mm) 4.58 6.51 2.5E+02 4.96 27.73 6.76 6.76
+RPE R.(deg.) 0.05 0.08 3.35 0.05 0.40 0.07 0.07
+Balloon2 ATE(mm) 7.81 13.51 3.5E+02 7.74 41.25 10.22 9.92
+RPE T.(mm) 12.82 14.16 3.9E+02 11.47 77.12 16.80 16.78
+RPE R.(deg.) 0.13 0.13 3.21 0.11 0.84 0.17 0.17
+DynamicFace ATE(mm) 2.80 9.71 1.1E+02 3.59 32.79 4.11 3.73
+RPE T.(mm) 1.71 7.93 2.4E+02 2.05 48.39 3.20 3.16
+RPE R.(deg.) 0.04 0.17 3.34 0.05 1.04 0.07 0.06
+Jumping ATE(mm) 7.65 13.31 2.8E+02 7.74 24.24 8.38 8.61
+RPE T.(mm) 10.25 11.27 2.8E+02 8.69 36.34 11.35 12.13
+RPE R.(deg.) 0.05 0.06 3.09 0.05 0.21 0.07 0.08
+Playground ATE(mm) 7.62 85.47 1.1E+02 5.45 27.44 6.47 5.06
+RPE T.(mm) 9.51 90.10 3.3E+02 7.68 40.28 8.00 6.42
+RPE R.(deg.) 0.10 0.75 4.69 0.11 0.40 0.10 0.10
+Skating ATE(mm) 7.21 19.37 78.24 7.28 27.57 9.21 8.88
+RPE T.(mm) 8.64 24.76 3.2E+02 8.65 45.02 11.19 11.44
+RPE R.(deg.) 0.04 0.15 3.91 0.05 0.24 0.07 0.07
+Truck ATE(mm) 22.55 - 1.1E+02 17.70 42.47 19.49 17.53
+RPE T.(mm) 31.68 - 3.6E+02 30.24 69.22 34.61 28.37
+RPE R.(deg.) 0.06 - 2.77 0.05 0.26 0.05 0.05
+Umbrella ATE(mm) 5.20 39.45 66.01 7.38 39.27 7.33 5.99
+RPE T.(mm) 8.11 12.08 3.7E+02 7.01 39.85 6.98 8.03
+RPE R.(deg.) 0.04 0.05 3.05 0.03 0.20 0.03 0.03
+Mean ATE(mm) 7.96 26.66 1.6E+02 7.81 32.06 8.67 7.98
+RPE T.(mm) 10.91 23.83 3.2E+02 10.09 47.99 12.36 11.64
+RPE R.(deg.) 0.07 0.20 3.43 0.06 0.45 0.08 0.08
+STD ATE(mm) 6.25 28.13 1.0E+02 4.25 8.11 4.89 4.50
+RPE T.(mm) 9.06 29.82 55.63 8.60 16.82 9.86 7.95
+RPE R.(deg.) 0.03 0.25 0.61 0.03 0.32 0.04 0.04